### v1

# Fuss-Free Network: A Simplified and Efficient Neural Network for Crowd Counting

Abstract
In the field of crowd-counting research, many recent deep learning based methods have demonstrated robust capabilities for accurately estimating crowd sizes. However, the enhancement in their performance often arises from an increase in the complexity of the model structure. This paper introduces the Fuss-Free Network (FFNet), a crowd counting deep learning model that is characterized by its simplicity and efficiency in terms of its structure. The model comprises only a backbone of a neural network and a multi-scale feature fusion structure. The multi-scale feature fusion structure is a simple architecture consisting of three branches, each only equipped with a focus transition module, and combines the features from these branches through the concatenation operation. Our proposed crowd counting model is trained and evaluated on four widely used public datasets, and it achieves accuracy that is comparable to that of existing complex models. The experimental results further indicate that excellent performance in crowd counting tasks can also be achieved by utilizing a simple, low-parameter, and computationally efficient neural network structure.
keywords:
crowd counting , Fuss-Free Network , multi-scale feature fusion , focus transition module , simplified structure
journal: Elsevier
[…]
[inst1]organization=School of Opto-Electronic and Communication Engineering, Xiamen University of Technology, city=Xiamen, country=China
1 Introduction
With the acceleration of urbanization and the increase in population mobility, effective monitoring and management of population density have become increasingly important. Rational regulation of population density not only concerns public safety but also affects social stability and the optimization of resource allocation [1]. Crowd counting technology, which has been receiving growing attention in the field of computer vision [2], has the potential to monitor and analyze crowd density in real-time, thereby helping to prevent safety incidents resulting from overcrowding. Moreover, it can provide essential data support for urban management, facilitating the optimization of resource allocation in crowded areas. Additionally, during the pandemic, this technology has the ability to monitor and regulate crowd movements, contribute to the development of public health strategies, and mitigate the transmission of the virus.
Figure 1: Representative images for crowd counting. These images illustrates the main challenges of crowd counting tasks, which include object occlusion and overlap in high density crowds, variations in scale and density within images, as well as the influence of lighting and background discrepancies, among other.
Crowd counting is a highly challenging computer vision task. Figure 1 shows three representative images for crowd counting, illustrating the following challenges in crowd counting tasks: (1) Variations in scale and density within an image: In general, crowds that are closer to the camera tend to have larger scales and lower densities, while those further away exhibit the opposite characteristics. (2) Object occlusion and overlap in high density crowds: High density crowd counting tasks frequently involve analyzing crowd images where a significant number of individuals are present, and situations such as object occlusion and overlap between individuals are commonly encountered. (3) The effects of lighting and background variations: The characteristics of individual features can vary due to differences in lighting conditions, which can include the creation of shadows. Moreover, the complexity and variability of the background environment can introduce interference factors, such as object occlusion. Moverover, crowd counting often involves the real-time processing of significant amounts of image or video data to fulfill the requirements of real-time crowd flow monitoring.
To overcome these challenges, mainstream methods of crowd counting rely on the powerful performance of deep learning models. These models are designed and trained to generate crowd density images at the pixel level. Subsequently, an accurate estimation of the total number of individuals is achieved by using a pixel-level crowd density image. To address the challenge of scale and density variations present in images, numerous methods have employed a multi-branch structure. This approach allows for the utilization of multi-scale features, effectively tackling the issue. Among them, some methods employ dilated convolutions with different dilation rates in a multi-branch structure, enabling the model to have receptive fields of different scales. Furthermore, to tackle challenges such as object occlusion, overlap in high density crowds, and the effects of lighting and background variations, almost all methods focus on enhancing the feature extraction capability of the model, primarily by selecting or designing various network architectures and employing different attention mechanisms. Optimizing the loss function yields significant effectiveness as well. For instance, in order to enhance model supervision, the loss function can be adjusted to ensure that the model establishes a more precise correspondence between the predicted density and the labeled location.
Different research works adopt different perspectives and consider various rules. Through intricate and sophisticated network structural designs and algorithm optimizations, existing deep learning models have made remarkable progress in improving the accuracy of crowd counting. However, the complexity of the model to a large extent hinders subsequent researchers from modifying and further improving the model. For example, efforts to simplify existing models with complex structure to reduce the consumption of computing resources may significantly impact the accuracy of crowd counting and the models’ adaptability in complex scenarios. However, for structurally simple models, it is relatively easier to modify the models while minimizing the performance decrease. Hence, the reduction in complexity of the model structure is beneficial for the subsequent advancement of crowd counting. In the pursuit of simple structure but also need to consider the counting performance, achieving a balance between simplifying structure and enhancing performance is a challenge that we need to pay attention to.
For the aforementioned reasons, our research objective is to design a crowd-counting model that strikes a balance between simplifying the model structure and improving counting performance for crowd counting. In terms of specific implementation, we first constructed a powerful backbone network, which is used to accurately capture the initial features of crowds from raw images. Next, we intricately designed a focus transition module, which efficiently integrates and refines key feature information in dynamic crowd scenes, ensuring the model’s effective adaptation to complex environmental changes. Finally, by implementing a practical and feasible multi-scale feature fusion strategy, we achieved deep integration and optimization processing of the extracted multi-level features. This ensures accurate identification and counting of crowds at different scales. Research shows that excellent crowd counting models can be designed with just a simple structure.
In summary, this paper’s contributions can be outlined in three aspects:
  1. 1.
    We propose a crowd-counting neural network with a simple structure that consists of only a backbone and a multi-scale feature fusion structure, in contrast to other complex models. However, it achieves results comparable to the state-of-the-art, even with low parameters and computational requirements.
  2. 2.
    We investigated three multi-scale feature fusion methods, including concatenate feature fusion, addition feature fusion, and stepwise addition feature fusion. Experimental results demonstrate that concatenate feature fusion has advantages in crowd-counting tasks. It efficiently extracts crowd features and effectively couples them across different scales, enhancing feature representation capability.
  3. 3.
    To effectively extract key information and reduce dimensionality, we have developed a focus transition module that integrates dynamic convolution. It can dynamically adjust multi-dimensional weights based on feature attributes and optimize weights in channel and spatial dimensions.
2 Related Work
Crowd counting is a task of estimating the number of people in a specific location. Early research employed human detection methods [3, 4], where researchers used detectors to detect human in images and accumulated the count of each detected person to obtain the crowd count result in the image. The limitation of this method lies in the significant impact of occlusions between persons in crowded environments or densely packed crowds on the performance of detectors, leading to a decrease in counting accuracy. Subsequently, direct regression methods [5] were proposed, which count the crowd directly from images. However, this method often cannot provide detailed information about individual distribution and positioning, and it is sensitive to model assumptions and parameter selection, which can affect accuracy in complex or highly variable density scenes. Therefore, methods based on density map regression [6, 7] have become mainstream, where the final crowd count is obtained by integrating density map estimates from images. Relying on the powerful representation capability of convolutional neural networks (CNNs), the majority of methods in this field are built upon the framework of CNNs. There are also many methods leveraging the strong feature extraction ability of transformers. To further improve performance, researchers have developed and integrated various technological and architectural innovations. In this section, we will review the relevant work on crowd counting from three aspects: model design, loss function design, and algorithm strategy design.
2.1 Model Design for Crowd Counting
With the development of deep learning, various methods for model design in crowd counting have emerged. These methods include introducing multi-scale fusion strategies, integrating attention mechanisms, and employing transformers. Multi-scale feature fusion methods attempt to address the challenge of large variations in crowd scale or size diversity by integrating features from multiple different levels. Zhang et al. [6] adopt a multi-column CNN architecture, where each column CNN utilizes filters with different receptive fields to adapt to scale variations. Jiang et al. [9] adopt an encoder-decoder structure that integrates multi-level features through skip connections across layers, aggregating low-level spatial details and high-level semantic information in a bottom-up manner. This enables accurate prediction from the entire image input to high-resolution density maps. Liu et al. [10] propose a novel deep learning architecture that utilizes spatial pyramid pooling to extract context information at different scales. It then uses these scale-aware features to regress the final density map. Ma et al. [11] introduce a multi-scale mechanism that utilizes information provided by different layers in the encoder to generate feature maps with varying receptive field sizes. Zhang et al. [12] designed a scaled adaptive convolutional neural network called SaCNN. This network combines features from multiple levels to enable the model to flexibly correspond to scale and viewpoint differences in the input image. Han et al. [13] employ a feature selection and inheritance learning approach, along with mask selection and inheritance loss, to perform multi-scale fusion. These methods achieve good results by effectively integrating multi-scale features.
Attention mechanisms are also an efficient method. Cao et al. [14] proposed SANet, which utilizes multi-scale convolutions and an adaptive global attention mechanism to handle crowd images of different sizes and densities. Jiang et al. [15] adopt a dual-branch structure, where one branch generates adjusted intermediate density maps and scale factors. It then corrects the unevenness of crowd density in the image by combining it with attention masks generated by the other branch, ultimately producing refined attention-based density maps. Lin et al. [16] employ a learnable global attention module and a local partition attention module, achieving decent performance in counting accuracy.
Given the remarkable success of transformer networks in the field of natural language processing (NLP), researchers have also begun to explore their application in crowd counting as a core feature extraction module. Sun et al. [17] utilize transformer models for feature extraction and information exchange by overlapping and concatenating image blocks and adding contextual tokens. Qian et al. [18] introduce a “U-shaped” multi-scale transformer network for crowd counting and introduce a new loss function to supervise regression, significantly improving counting accuracy.
In addition to the aforementioned methods, there are some other designs that contribute to crowd counting. The CSRNet proposed by Yan et al. [2] consists of a convolutional neural network serving as the frontend for 2D feature extraction, and it utilizes dilated convolutional neural networks as the backend to provide a larger receptive field and replace pooling operations. Similarly, Tian et al. [8], besides employing a transformer network that focuses more on spatial aspects, also utilize a regression head composed of dilated convolutions to expand the receptive field, achieving good counting results as well.
2.2 Loss function design for crowd counting
Many studies have designed loss functions tailored to the characteristics of crowd counting. Ma et al. [20] no longer rely on low-quality density maps generated from Gaussian kernels but construct a density contribution probability model from point annotations, using Bayesian inference for loss estimation. Wang et al. [21] treat crowd counting as a distribution matching problem, employing Optimal Transport (OT) to find the best match between the generated density map and the point annotations. Yan et al. [22] proposed a multi-resolution loss function, PML, which can learn to better fit the posterior distribution using multi-scale information, thereby improving counting performance.
2.3 Algorithmic Strategies design for crowd counting
In addition to model design and loss function design, there are also studies that approach crowd counting from algorithmic perspectives. Song et al. [23] developed a novel approach named P2PNet, which directly predicts head coordinates and their confidences in images and matches points to their ground-truth targets using the Hungarian algorithm. Lin et al. [19] decompose the pedestrian counting process into point-querying processes and use a custom quad-tree structure to adapt to crowds of different densities. Additionally, it employs a recursive rectangular window attention mechanism to improve computational efficiency.
3 Method
Figure 2: This figure presents the architectural designs of four different crowd counting models, including SASNet, STEERER, CCTrans, and our simplified architecture.
With the continuous evolution and development of deep learning technology, high-performance crowd counting models have achieved significant improvements in accuracy, but they also face the problem of increasing model complexity. As depicted in Figure 2, the three exemplary high-performance models (CCTrans, STEERER, and SASNet) demonstrate exceptional accuracy, but their underlying structures are highly complex. Specifically, SASNet utilizes multiple decoder blocks along with a complex fusion header design and interleaved fusion mechanisms, resulting in a complex architecture. STEERER relies on the intricate and computationally demanding HRNet48 as its backbone network and designs a complex fusion module. Although the CCTrans appears to be lightweight, it actually adds more complexity to the model overall because it is based on the massive parameter Twins-SVT pre-training network, even though it only attaches a regression header after multi-branch fusion.
Based on this, the core of our research is to conceive and build a more streamlined and efficient framework for crowd counting models. The fourth framework in Figure 2 is precisely the simplified solution we propose. This model adopts the compact yet powerful ConvNeXT-Tiny as the base backbone network, and it comprises a multi-scale feature fusion structure that includes a focus transition module. We refer to this model as Fuss-Free Network(FFNet).
FFNet is specifically illustrated in Figure 3. Firstly, the input image goes through the feature extraction backbone in the framework to extract features at three different scales. Then, these three features pass through the focus transition module to generate more semantically rich features. Next, each stage’s features go through channel concatenation and spatially-aware fusion via the framework’s multi-scale fusion. Finally, regression heads are used to regress the density maps. The final density maps and the integrated counts will be used to construct the loss function.
Figure 3: The network architecture of FFNet. H, W, C denote the height, width, and number of channels of the features, respectively. Gap represents global average pooling, Mean denotes mean operation, Conv stands for convolution, $\times1$Conv denotes convolution with a $\times1$ kernel, Dyconv represents dynamic convolution, MLP refers to multilayer perceptron, Sigmoid is the activation function, and the circles with “×”, “+”, and “C” represent element-wise multiplication, element-wise addition, and channel concatenation operations, respectively.
3.1 Feature Extraction Backbone
Different from most crowd counting studies, we employ ConvNeXt-Tiny [24] as the feature extraction backbone for the multi-scale fusion framework. This network integrates key design principles of transformers, such as layer normalization and multi-head attention mechanisms, and implements them in the ConvNeXt block structure. This combination allows ConvNeXt-Tiny to not only maintain the powerful processing capability of CNNs for spatial local features but also significantly enhance the understanding and utilization of global information.
Compared to the deep structures of VGG [25] and the residual learning of ResNet [26], ConvNeXt-Tiny provides a more efficient computational path and a simplified network architecture. Its streamlined and efficient characteristics ensure that the model can achieve superior performance in visual tasks such as image recognition compared to traditional models. Therefore, it also aligns better with our desire for a concise and efficient network design. For crowd counting tasks, dealing with objects of various scales has always been a significant challenge. A widely used solution is to generate features at different resolutions. Thus, we will utilize the ConvNeXt-Tiny pretrained model to build a multi-stage hierarchical feature extraction structure, where each stage can capture feature representations at different levels and resolutions. This approach enables us to better address the multi-scale challenges of crowd counting while ensuring the simplicity of the framework.
3.2 Multi-scale Fusion Strategy
Figure 4: The three simple designs for multi-scale fusion are illustrated below. Given $S_{1}$, $S_{2}$, and $S_{3}$ as inputs, T represents the final output, and $T_{2}$ represents the output of the $S_{2}$ stage.
For features at different levels of resolution, we need to align and fuse them. To avoid introducing too much additional computational cost and adhere to the principle of simplicity and efficiency, we propose three simple designs for multi-scale fusion. As shown in the figure 3, the “Concatenate” fusion directly applies transposed convolution to features $S_{3}$ and $S_{2}$ for upsampling to match the spatial size of $S_{1}$, and then fuses the features through channel concatenation. The “Addition” fusion first reduces the number of channels in features $S_{3}$ and $S_{2}$ to match the channel count of $S_{1}$. Before performing element-wise addition, transposed convolution operations are used to align the spatial size of the features. The “Stepwise Addition” fusion involves progressively upsampling and combining features of different scales, similar to the design of FPN [28]. In detail, these designs employ transpose convolution instead of interpolation operations because transpose convolution can better preserve and reconstruct the spatial structure information and feature hierarchy of the image. Compared to simple methods like bilinear interpolation for upsampling, transpose convolution enhances the output resolution by learning weight parameters, allowing for a more flexible and adaptive generation of high-quality feature maps. This is particularly significant for reconstructing complex crowd distribution scenes. The above three designs performed well in experiments, with the “Concatenate” fusion achieving the best results. We default to this design in the FFNet framework.
3.3 Focus Transition Module
Directly fusing features from different levels and resolutions of the backbone may lead to information redundancy due to misalignment in spatial dimensions and an excessive number of channels. This could potentially result in adverse effects such as decreased learning efficiency and inaccurate information propagation. Especially in crowd counting tasks, directly fusing features may fail to accurately retain spatial details from lower layers and lack an understanding of the global scene layout. This can lead to biases in estimating dense or sparse crowd distributions. Therefore, to address this issue and fully exploit the potential of the framework, we designed a focus transition module. It effectively focuses on dynamic and static features, achieving efficient dimensionality reduction and feature extraction to facilitate the transition.
Inspired by [27], we employed a novel dynamic convolution in the design stage, as illustrated in Figure 5.
Figure 5: Dyconv. FC stands for fully connected layer, Gap represents global average pooling, Conv denotes convolution, Sigmoid and Relu are activation functions, multiply refers to element-wise multiplication, $A_{nk}$, $A_{sk}$, $A_{ik}$, $A_{ok}$ represent attention along the dimensions of the number of convolution kernels, convolution kernel size, input channels, and output channels, respectively, and $W_{k}$ represents the weights of different convolution kernels.
Its core idea is to dynamically generate convolution kernel weights during the model runtime based on the input data and the network’s learned state, rather than using predefined and fixed weight matrices. Dynamic convolution will flexibly capture features across the four key dimensions: input channels, output channels, kernel size, and the number of kernels within the convolution. It utilizes average pooling and convolution for attention and then dynamically weights them. It can adaptively adjust the parameters of the convolutional kernel to better match the features of the input data. This adaptive capability enables dynamic convolution to effectively extract features from data of different scales and complexities, thereby enhancing the model’s ability to handle diverse inputs. In terms of algorithms, we assume X to be the input of dynamic convolution. The process of dynamic convolution attention extraction and dynamic weighting can be defined as:
$A(\mathbf{X})$ $=\sigma(\operatorname{Conv}(FC(\operatorname{AvgPool}(\textbf{X})))),$ (1)
$Y(\mathbf{X})$ $=\mathbf{X}\odot\sum_{k=1}^{n}(\mathbf{A_{nk}}\odot\mathbf{A_{sk}}\odot\mathbf{A_{ik}}\odot\mathbf{A_{ok}}\odot\mathbf{W_{k}}),$ (2)
where A(X) represents the attention calculation, AvgPool denotes average pooling, FC stands for fully connected layer, Conv indicates convolution operation, n represents the number of convolutional kernels, and $\sigma$ is the sigmoid activation function; Y(X) represents the dynamic convolution weighting process, $A_{nk}$, $A_{sk}$, $A_{ik}$, $A_{ok}$ are the attention weights on four dimensions: the number of convolution kernels, kernel size, input channels, and output channels, respectively, $W_{k}$ represents the weights of different convolution kernels, and C denotes the input optimized by weights.
In terms of specific design, as illustrated in Figure 3, the entire structure combines feature weight optimization, dynamic perception, and residual components. The feature output from the backbone is first subjected to adaptive feature weight optimization. This mechanism strengthens and filters features along the channel dimension, ensuring that the model focuses on the features crucial for crowd statistics. Then, the stacked dynamic convolutions automatically adjust filter parameters for different inputs to capture dynamic and static feature changes in the scene more effectively, achieving feature extraction in the process. Finally, the module performs spatial feature weight optimization, allocating different attention to each pixel in the image, enabling the model to automatically focus on key areas of crowd distribution while ignoring irrelevant background noise. In addition, we embed residual structures throughout the entire process to ensure that the original low-level feature details are preserved and integrated with the higher-level abstract features refined through multiple convolutional layers. It is worth noting that while the focus transition module enhances key features, it also implicitly achieves feature dimensionality reduction. This dimensionality compression and optimization, while preserving the integrity of key information, help reduce the computational complexity of the model, thereby improving operational efficiency. In detail, after each dynamic convolution, there is a ReLU activation layer and a BatchNorm layer to ensure the non-linear expression capability of features and the stability of normalization. The entire process of the focus transition can be represented by the following formula:
$C(\mathbf{F})$ $=\sigma(\operatorname{MLP}(\operatorname{AvgPool}(\mathbf{F})))\odot\mathbf{F},$ (3)
$P(\mathbf{C})$ $=Y[Y(Y(Y(\mathbf{C})+\mathbf{C})+(Y(\mathbf{C})+\mathbf{C}))],$ (4)
$S(\mathbf{P})$ $=\sigma(\operatorname{Conv}(\mathbf{\bar{P}}))\odot\mathbf{P},$ (5)
where F is the input, C(·) is the channel feature weight optimization, Y(·) is the dynamic convolution, P(·) represents the output after stacked dynamic convolutions, S(·) is the spatial feature weight optimization, $\bar{P}$ is the mean value of P, and MLP is the multi-layer perceptron.
The focus transition module we designed will fully unleash the potential of the streamlined framework, aiding the model in extracting richer, more targeted semantic information, achieving smooth transition effects, and enhancing the accuracy and robustness of crowd counting tasks.
3.4 Loss Function
The total loss consists of three components: count loss, optimal transport(OT) loss [20], and variation loss.
Count loss is a critical loss in crowd counting that quantifies the gap between the number of people predicted by the model and the actual number of people labeled, thus contributing to better convergence of the model. It is defined as the formula 6. Based on the counting loss, we further constructed a variation loss using the predicted density map and the ground truth density map. It employs the L1 norm to measure the intensity or grayscale variation differences between adjacent pixels in the predicted and ground truth density maps. Through this approach, our goal is to enhance the smoothness of the image and effectively reduce the adverse effects of potential noise on density estimation. Its definition is as follows:
$\mathcal{L}_{c}$ $=\frac{1}{N}\sum_{i=1}^{N}\left|c_{p}-c_{g}\right|,\mathcal{L}_{v}=\frac{1}{N}\sum_{i=1}^{N}\left|x_{i}-y_{i}\right|,$ (6)
where $c_{p}$ is the predicted population and $c_{g}$ is the actual population, N represents the batch size, $x_{i}$ represents the density value of the $i$-th pixel in the true density map, and $y_{i}$ represents the predicted density value of the $i$-th pixel at the corresponding position in the predicted density map.
OTLoss is used to measure the difference between two probability distributions, making it particularly suitable for handling misaligned data distribution problems. It encourages the model to learn probability densities that are closer to the true distribution. It first calculates the distance matrix between the predicted density map and the true density map using the Euclidean distance. Then, it utilizes the Sinkhorn algorithm for iterative optimization, finding the optimal transportation plan from the source distribution to the target distribution, and obtaining the dual variables during this process. Finally, based on these dual variables and the predicted density distribution, OTLoss is computed. This loss measures the discrepancy between the predicted density distribution and the true labels in terms of optimal transportation. Throughout the entire process, the regularization parameter regulates the balance between distance cost and distribution smoothness. Its complete definition is as follows:
$D_{ij}$ $=\sqrt{\left(x_{i}-x_{j}\right)^{2}+\left(y_{i}-y_{j}\right)^{2}},$ (7)
where each matrix element $D_{ij}$ represents the Euclidean distance between the $i$-th grid point in the predicted density map, with coordinates ($x_{i}$, $y_{i}$), and the $j$-th ground-truth point in the true density map, with coordinates ($x_{i}$, $y_{j}$).
$\mathbf{\beta}=\log(\mathbf{P})-\frac{1}{\lambda}\left[\log\left(\sum_{j=1}^{N}e^{-\lambda D_{ij}}\right)\right]_{i=1}^{M},$ (8)
where $\beta$ represents the dual variables obtained through the Sinkhorn algorithm, $D$ is the distance matrix, $P$ represents the prediction map, $M$ represents the number of rows in matrix $D$, NN represents the number of columns, and $D_{ij}$ represents the value at the $i$-th row and $j$-th column of the distance matrix, and $\lambda$ is the regularization parameter set to 10.
$\mathcal{L}_{ot}$ $=\sum_{i=1}^{N}\left[\left(\frac{\mathbf{\beta}_{i}}{P_{i}}-\frac{\mathbf{\beta}_{i}\cdot\mathbf{\rho}_{i}}{P_{i}^{2}}\right)\mathbf{\rho}_{i}\right],$ (9)
where $N$ represents the batch size, $\beta_{i}$ is the dual variable computed by the Sinkhorn algorithm for the $i$-th sample, $P_{i}$ is the sum of predicted densities in the $i$-th sample, and $\rho_{i}$ is the vector composed of all predicted density values for the $i$-th sample.
Finally, the total loss is the weighted sum of three components. Let $\lambda_{1}$ and $\lambda_{2}$ be the weighting values, the total loss is defined as:
$\mathcal{L}_{all}$ $=\mathcal{L}_{c}+\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{2},$ (10)
Combining counting loss, OTLoss, and variation loss can enhance the accuracy and robustness of crowd counting models. Additionally, regularization reduces noise in the predicted maps, improves model generalization, accelerates convergence, and helps train more accurate and stable crowd counting models. In our experiments, we fixed $\lambda_{1}$ at 0.1 and only fine-tuned $\lambda_{2}$.
4 Experiment
This section validates the effectiveness of our method on crowd counting benchmarks through extensive experiments and ablation studies, comparing it with state-of-the-art techniques.
4.1 Implementation Details
4.1.1 Dataset
We evaluate our methods in four benchmarks, including UCF_CC_50 [29], ShanghaiTech Part A and B [6], and NWPU-Crowd [30]. These datasets differ in terms of scene complexity, crowd density range, and image resolution.
UCF_CC_50 [29] consists of 50 images of high-density crowds, with a total of 63,075 individuals across the entire dataset. The number of individuals in each image ranges from 94 to 4543, with an average of 1280 individuals per image. Extreme variations in crowd density and occlusion situations are among the challenges presented by the densely crowded complex scenes in this dataset. Due to the relatively small size of the dataset, Idrees et al. [29] defined a cross-validation protocol for training and validating models.
ShanghaiTech [6] comprises a total of 1198 annotated images with 330,165 individually marked heads, divided into Part A and Part B. Part A consists of 482 images, among which 300 are allocated for training and 182 for testing; these images were sourced from the Internet. In contrast, Part B comprises 400 training images and 316 testing images, all sourced from real surveillance footage. These two parts of the dataset, due to their complex real-life scenarios, diverse environmental conditions, and distribution of crowds, pose significant challenges for crowd counting technology. They are crucial benchmarks for measuring and improving the accuracy of crowd counting algorithms.
NWPU-Crowd [30] differs from the previous ones in several aspects. It has a larger scale, containing 5,109 high-resolution images covering various lighting conditions, perspectives, and scene types. The number of people ranges from tens to tens of thousands, providing more diverse scenes and a wider range of people counts. Additionally, the dataset includes a few negative samples that contribute to the model’s increased robustness. Compared to the previous datasets, it imposes higher precision requirements and tests the generalization ability of crowd counting algorithms.
4.1.2 Training Settings and Hyperparameters
For a fair comparison, the backbone of FFNet is the official ConvNeXt-Tiny model, which is pretrained on the ImageNet-1k dataset. In our experiments, data augmentation was limited to random cropping and random horizontal flipping. For ShanghaiTech Part B, we set the cropping size to 512; for NWPU, the cropping size is set to 384; and for UCF_CC_50 and ShanghaiTech Part A datasets, we adjust the cropping size to 256. The model training utilized the AdamW optimizer with a batch size of 8, and the initial learning rate was set to 1e-5. Additionally, to mitigate overfitting, we introduced an L2 regularization term with a coefficient of 0.005. All experiments were conducted using PyTorch on a single 16-GB RTX 4080.
4.1.3 Evaluation Metrics
In crowd counting research, Mean Absolute Error (MAE) and Mean Squared Error (MSE) are commonly used as evaluation metrics for crowd counting. The Mean Absolute Error (MAE) is the average of the absolute differences between the predicted counts and the actual counts. It provides an intuitive measure of how much the predicted values deviate from the actual values. The Mean Squared Error (MSE) is the average of the squared differences between the predicted counts and the actual counts. It is highly sensitive to outliers and requires the model to have good robustness to noise and outliers in the data. The expressions for MAE and MSE are as follows:
$MAE$ $=\frac{1}{N}\sum_{i=1}^{N}\left|c_{i}-\hat{c}_{i}\right|,$ (11)
$MSE$ $=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(c_{i}-\hat{c}_{i}\right)^{2}},$ (12)
where N is the number of test images, and $c_{i}$ and $\hat{c}_{i}$ are the predicted and true number of people in the ith image, respectively.
Method Venue Backbone Parameters FLOPS UCF_CC_50 SHA SHB NWPU(V) MAE MSE MAE MSE MAE MSE MAE MSE MCNN [6] CVPR16 CNN - - 377.6 509.1 110.2 173.2 26.4 41.3 218.5 700.6 CSRNet [2] CVPR18 VGG-16 - - 266.1 397.5 68.2 115.0 10.6 16.0 104.9 433.5 CAN [10] CVPR19 VGG-16 - - 212.2 243.7 62.3 100.0 7.8 12.2 93.5 489.9 SFCN [31] CVPR19 ResNet-101 - - 214.2 318.2 64.8 107.5 7.6 13.0 - - BL [20] ICCV19 VGG-19 - - 229.3 308.2 62.8 101.8 7.7 12.7 93.6 470.4 ASNet [15] CVPR20 VGG-16 - - 174.8 251.6 57.8 90.1 - - - - AMRNet [32] ECCV20 VGG-16 - - 184.0 265.8 61.5 98.3 7.0 11.0 - - DM-Count [21] NeurIPS20 VGG-19 - - 211.0 291.5 59.7 95.7 7.4 11.8 70.5 357.6 P2PNet [23] ICCV21 VGG-16 19.22M 104.71G 172.7 256.2 52.7 85.1 6.3 9.9 - - DKPNet [33] ICCV21 HRNet-W40 12.60M 46.31G - - 55.6 91.0 6.6 10.9 61.8 438.7 SASNet [38] AAAI21 VGG-16 38.90M 232.98G 161.4 234.46 53.5 88.3 6.3 9.9 - - MAN [16] CVPR22 VGG-19 30.96M 105.85G - - 56.8 90.3 - - - - CHFL [34] CVPR22 VGG-19 21.51M 108.27G - - 57.5 94.3 6.9 11.0 76.8 343.0 GauNet [35] CVPR22 ResNet-50 - - 186.3 256.5 54.8 89.1 6.2 9.9 - - CLTR [36] ECCV22 ResNet-50 40.95M 28.73G - - 56.9 95.2 6.5 10.6 61.9 246.3 FIDT [40] TMM22 HRNet-W48 66.58M 142.40G 171.4 233.1 57.0 103.4 6.9 11.8 51.4 107.6 PML [22] TCSVT23 ConvNeXt-S - - - - 53.4 93.1 6.2 9.7 - - LDFNet [37] arXiv23 VGG-19 - - 164.9 234.8 52.7 82.9 6.7 11.5 - - PET [19] ICCV23 VGG-16 51.65M 859.68G - - 49.3 78.9 6.2 9.7 58.5 238.0 STEERER [13] ICCV23 HRNet-W48 64.42M 94.40G - - 54.5 86.9 5.8 8.5 54.3 238.3 BCCT [17] arXiv21 T2T-ViT-14 - - - - 53.1 82.2 7.3 11.3 53.0 170.3 CCTrans [8] arXiv21 Twins-SVT 103.58M 99.44B 168.7 234.5 52.3 84.9 6.2 9.9 38.6 87.8 CUT [18] BMVC22 Twins-PCPVT 68.09M 90.45G - - 51.9 79.1 - - - - FFNet(ours) - ConvNeXt-T 29.02M 23.67G 161.1 233.3 48.3 80.5 6.1 9.0 41.2 113.2
Table 1: Compared with the advanced methods on UCF_CC_50, ShanghaiTech Part A and B, and NWPU datasets. And compare the parameter quantity and flops with advanced methods. The best results are shown in bold, and the second-best results are underlined.
4.2 Comparison with state-of-the-art methods
4.2.1 Counting performance
CNN-based crowd counting methods dominate the field of crowd counting, with most approaches building upon the foundation of CNNs for model design. As shown in Table 1, our method outperforms them, exhibiting significant advantages in several benchmark tests and achieving state-of-the-art (SOTA) or near-state-of-the-art performance levels. Even when compared to the popular transformer-based methods in recent years, our model demonstrates superiority, achieving better results in most cases.
Specifically, we achieved outstanding results on the UCF_CC_50 dataset, with an MAE of 161.4, surpassing existing methods and reaching a new state-of-the-art level. On the ShanghaiTech Part A subset, we demonstrated significant superiority with an MAE value of 48.3, successfully surpassing the latest SOTA in this field. In the evaluation of the ShanghaiTech Part B subset, despite fierce competition, our method still achieved a highly competitive second-place performance with an MAE of 6.1. Finally, in the more challenging NWPU-Crowd dataset, our method also achieved results close to the top tier. Without excessive complexity in design, our model achieves incredible results on these datasets. This fully demonstrates the effectiveness of FFNet.
4.2.2 Parameters and Computational Resources
Adhering to the original intention of our work, we aimed to simplify the model while achieving high crowd counting performance. Therefore, we conducted lightweight testing on FFNet and state-of-the-art methods. The testing primarily focused on open-source models released in the past three years. We did not test some methods that are partially open-source or not open-source. For a fair comparison, we all test with an all-one tensor of size (1, 3, 512, 512) as input. And we measured the lightweight characteristics of the model based on two metrics: model parameters and floating-point operations (FLOPs). Among them, the model parameters is one of the key indicators for measuring the complexity of neural network models. It refers to the total number of weights and bias parameters in the model. Models with fewer parameters occupy less storage resources, have faster inference speeds, are easier to deploy, and have lower risks of overfitting. FLOPS is another important metric for measuring the complexity and computational requirements of neural network models. In the experiment, it refers to the number of floating-point operations required for one forward pass. Models with lower FLOPS consume fewer computational resources, have faster inference speeds, are more suitable for low-power environments, and are easier to deploy on mobile and embedded devices.
The experimental testing results are presented in Table 1. The data shows that over the past three years, as counting performance has continued to improve, there has been a significant increase in model parameters and FLOPs, resulting in greater model complexity. However, FFNet’s parameters and FLOPs are significantly lower than those of other advanced crowd counting methods. While achieving superior counting performance, achieve optimal FLOPs and highly competitive model parameters. We believe that FFNet’s SOTA counting performance with such low FLOPs and model parameters is truly commendable. This effectively proves the efficacy of our proposed model: concise, efficient, low in complexity, and yet powerful in performance. FFNet effectively balances model simplification and performance enhancement, making it suitable not only for deployment in resource-constrained environments but also for performing exceptionally well in scenarios with high real-time requirements. This significantly drives the development and application of efficient crowd-counting technologies.
4.3 The Effectiveness of Our Simplified Structure
To validate the effectiveness of our concise network framework, we conducted corresponding experiments. We replaced FTM with ordinary convolutions. Specifically, we replaced FTM with four concatenated simple convolutions on top of a simple framework.
Method SHA SHB
MAE MSE MAE MSE
SASNet [38] 53.5 88.3 6.3 9.9
CCTrans [8] 52.3 84.9 6.2 9.9
CUT [18] 51.9 79.1 - -
PET [19] 49.3 78.9 6.2 9.7
STEERER [13] 54.5 86.9 5.8 8.5
FFNet+CNN 51.7 89.0 6.2 9.5
FFNet+FTM 48.3 80.5 6.1 9.0
Table 2: Simple network experiment: A comparison between a simple network with ordinary convolutions and a simple network with FTM against state-of-the-art methods.
The results, as shown in Table 2, demonstrate that, compared to state-of-the-art approaches, our framework achieved an MAE of 51.7 on the ShanghaiTech Part A dataset and 6.2 on the Part B dataset, solely utilizing ordinary convolutions for intermediate downsampling and feature extraction. This surpasses most state-of-the-art methods. This further confirms the effectiveness of the streamlined network architecture we adopted and verifies the positive role of the FTM module in enhancing the overall network performance.
4.4 Multi-scale fusion experiment
In this section, we conducted comparative experiments on three simple multi-scale fusion strategies designed in Section 3.2. The experiment only adjusted the fusion method of the FFNet and validated it on the Part A and Part B datasets. The experimental results are shown in Table 3 below. In Part A, both the “Add” and “Stepwise addition” methods demonstrate good performance, but the “concatenate” method outperforms them. In Part B, the “concatenate” method shows better performance and robustness. This proves the effectiveness of the “concatenate” method in crowd counting tasks.
Fusion Strategy SHA SHB
MAE MSE MAE MSE
Add 52.6 87.5 6.4 10.8
Stepwise addition 52.3 91.9 7.0 11.7
concatenate 48.3 80.5 6.1 9.0
Table 3: Results of three simple multi-scale fusion designs on ShanghaiTech Part A and B datasets.
4.5 Ablation Study
As mentioned earlier, our proposed FFNet consists of the following components: a feature extraction backbone, focus transition modules, and a multi-scale fusion structure. We conducted extensive ablation experiments on the ShanghaiTechA dataset to validate the effectiveness of each component in FFNet.
4.5.1 Focus Transition Module
We conducted two ablation experiments on the FTM module, one focusing on local effects and the other on global effects, to verify its effectiveness. First, at the local level, we tested the counting results of the three branches of the model before simple fusion on dataset A. Then, based on this, we removed the FTM module and tested the counting results of the three branches again, and compared the results. The effectiveness of FTM will demonstrate by observing the difference of counting performance before and after the three branches. At the global level, we tested the performance of FFNet with and without FTM to highlight the overall contribution of FTM to FFNet. The results are shown in Table 4. The results indicate that employing the FTM module, whether in local branches or in the overall model, effectively improves the counting accuracy, demonstrating the practicality of the focus transition module.
Method SHA
MAE MSE
[…]
Table 4: Effective improvement of model counting performance by FTM.
4.5.2 Loss Function
We also quantitatively evaluated the weighted sum composed of counting loss, variation loss, and optimal transport loss to verify the specific contributions of each loss function to the model’s performance. The results are shown in Table 5. Experimental results show that the model performance reaches its optimal state when the $\lambda_{1}$ and $\lambda_{2}$ are 0.1 and 0.01, respectively. This weighting scheme highlights the central role of count loss in crowd counting prediction, while also moderately considering the image smoothness brought by variation loss and the fine depiction of complex scenes by OT Loss.
Loss function SHA
MAE MSE
$L_{c}$ 55.1 95.3
$L_{c}+L_{ot}+L_{v}$ 55.3 94.9
$L_{c}+\lambda_{1}L_{ot}+\lambda_{2}L_{v}(\lambda_{2}=0.1)$ 54.6 90.4
$L_{c}+\lambda_{1}L_{ot}+\lambda_{2}L_{v}(\lambda_{2}=0.01)$ 48.3 80.5
Table 5: The total loss function achieved optimal performance when weighting three loss functions and setting $\lambda 1$ and $\lambda 2$ to 0.1 and 0.01, respectively.
4.5.3 Composition Analysis of FFNet
Table 6 displays the component analysis of our model. For the ShanghaiTech PartA dataset, both FTM and the final “Concatenate” fusion contribute significantly to the final MAE metric.
Method SHA
MAE MSE
backbone 54.5 93.1
backbone+FTM 50.9 84.4
backbone+FTM+Concatenate 48.3 80.5
Table 6: Composition analysis of FFNet.
4.6 Visualization
In order to further explore the internal mechanism of the model, we visualized the effective receptive field maps and their heat maps of the three branches of the model before simple fusion, so as to reveal the spatial coverage and response intensity distribution of the input data features of each branch of the model before simple fusion. The effective receptive field [39] refers to a pixel on the output feature map of a certain layer in a neural network whose response is influenced by the size of the input image, reflecting the actual sampling range of the layer on the input spatial size. The heatmap is a visualization tool that represents numerical intensity through color coding and is used to display the model’s level of attention or confidence distribution towards specific targets or features in the input data.
As shown in Figure 6, the first row is the effective receptive field of three different branches, the second row is the corresponding heat map and e ach column represents a different branch of the FFNet. From the effective receptive field maps, we can see that the effective receptive field of different branches is different, and the deeper the branch, the larger the effective receptive field, which indicates that our model can capture information at different scales through three branches. The heatmap shows the degree of attention of each branch to the input image through the change in color shade. The first branch is more concerned with the low-level, local, and basic visual information of the image, such as edge, texture, color, etc., and has good maintenance of spatial details. In the second and third branches, attention gradually increases, focusing on abstract and macroscopic features. The color distribution of the heatmaps also becomes more concentrated, demonstrating the model’s ability to integrate global information and recognize the overall distribution and density of the crowd.
Figure 6: Three branches of the model pass FTM module in the receptive field map and their heat map
Finally, we visualize the model outputs using six images in the NWPU dataset, which exhibit varying scales, densities, lighting conditions and from different scenes. The visualization of the model’s final outputs is presented in Figure 7. Our model performs excellently on images of various scales, demonstrating good technical accuracy.
Figure 7: FFNet visualizations on the most challenging NWPR populations. Six images from different scenes, with different scales, densities, and lighting differences, show the counting and fitting capabilities of FFNet.
5 Conclusion
In this paper, we propose a concise and efficient crowd counting model called Fuss-Free Net work (FFNet), which consists only of a backbone network, a multi-scale fusion structure, and focus transition modules. In the concrete implementation, we use the ConvNeXt-Tiny network to build a multi-branch structure, each branch is equipped with a focus transition module, and then the multi-branch features are fused through a simple cascading fusion strategy. Finally, experiments show that the counting accuracy of the model is equal to or even better than that of the complex model, despite its compact structure, small number of parameters, and low computational complexity, breaking the inherent relationship between complexity and performance improvement.
FFNet provides a simple and efficient solution for crowd-counting tasks, but there is still plenty of room to explore. For example, cross-domain adaptability, interpretability, and robustness are improved. In our future work, we will explore the field of crowd counting from the perspectives of simplicity, robustness, and high counting accuracy, hoping to build a more general, reliable, and robust model.

### v3

@@
-Fuss-Free Network: A Simplified and Efficient Neural Network for Crowd Counting
+The Effectiveness of a Simplified Model Structure for Crowd Counting
 Abstract
-In the field of crowd-counting research, many recent deep learning based methods have demonstrated robust capabilities for accurately estimating crowd sizes.
+In the field of crowd counting research, many recent deep learning based methods have demonstrated robust capabilities for accurately estimating crowd sizes.
 However, the enhancement in their performance often arises from an increase in the complexity of the model structure.
-This paper introduces the Fuss-Free Network (FFNet), a crowd counting deep learning model that is characterized by its simplicity and efficiency in terms of its structure.
-The model comprises only a backbone of a neural network and a multi-scale feature fusion structure.
-The multi-scale feature fusion structure is a simple architecture consisting of three branches, each only equipped with a focus transition module, and combines the features from these branches through the concatenation operation.
+This paper discusses how to construct high-performance crowd counting models using only simple structures.
+We proposes the Fuss-Free Network (FFNet) that is characterized by its simple and efficieny structure, consisting of only a backbone network and a multi-scale feature fusion structure.
+The multi-scale feature fusion structure is a simple structure consisting of three branches, each only equipped with a focus transition module, and combines the features from these branches through the concatenation operation.
 Our proposed crowd counting model is trained and evaluated on four widely used public datasets, and it achieves accuracy that is comparable to that of existing complex models.
-The experimental results further indicate that excellent performance in crowd counting tasks can also be achieved by utilizing a simple, low-parameter, and computationally efficient neural network structure.
-keywords:
-crowd counting , Fuss-Free Network , multi-scale feature fusion , focus transition module , simplified structure
-journal: Elsevier
-[…]
-[inst1]organization=School of Opto-Electronic and Communication Engineering, Xiamen University of Technology, city=Xiamen, country=China
-1 Introduction
-With the acceleration of urbanization and the increase in population mobility, effective monitoring and management of population density have become increasingly important.
-Rational regulation of population density not only concerns public safety but also affects social stability and the optimization of resource allocation [1].
-Crowd counting technology, which has been receiving growing attention in the field of computer vision [2], has the potential to monitor and analyze crowd density in real-time, thereby helping to prevent safety incidents resulting from overcrowding.
+Furthermore, we conduct a comprehensive evaluation by replacing the existing backbones of various models such as FFNet and CCTrans with different networks, including MobileNet-v3, ConvNeXt-Tiny, and Swin-Transformer-Small.
+The experimental results further indicate that excellent crowd counting performance can be achieved with the simplied structure proposed by us.
+Index Terms:
+crowd counting, simplified structure, Fuss-Free Network, multi-scale feature fusion, focus transition module.
+I Introduction
+Crowd counting, which has been receiving growing attention in the field of computer vision [1], has the potential to monitor and analyze crowd density in real-time, thereby helping to prevent safety incidents resulting from overcrowding.
 Moreover, it can provide essential data support for urban management, facilitating the optimization of resource allocation in crowded areas.
-Additionally, during the pandemic, this technology has the ability to monitor and regulate crowd movements, contribute to the development of public health strategies, and mitigate the transmission of the virus.
+Additionally, during the pandemic, it has the ability to monitor and regulate crowd movements, contribute to the development of public health strategies, and mitigate the transmission of the virus.
 Figure 1: Representative images for crowd counting.
@@
 Figure 1 shows three representative images for crowd counting, illustrating the following challenges in crowd counting tasks: (1) Variations in scale and density within an image: In general, crowds that are closer to the camera tend to have larger scales and lower densities, while those further away exhibit the opposite characteristics.
-(2) Object occlusion and overlap in high density crowds: High density crowd counting tasks frequently involve analyzing crowd images where a significant number of individuals are present, and situations such as object occlusion and overlap between individuals are commonly encountered.
-(3) The effects of lighting and background variations: The characteristics of individual features can vary due to differences in lighting conditions, which can include the creation of shadows.
+(2) Overlap occlusion and object in high density crowds: High density crowd counting tasks frequently involve analyzing crowd images where a significant number of individuals are present, and situations such as object occlusion and overlap between individuals are commonly encountered.
+(3) Lighting and background variations: The characteristics of individual features can vary due to differences in lighting conditions, which can include the creation of shadows.
 Moreover, the complexity and variability of the background environment can introduce interference factors, such as object occlusion.
-Moverover, crowd counting often involves the real-time processing of significant amounts of image or video data to fulfill the requirements of real-time crowd flow monitoring.
 To overcome these challenges, mainstream methods of crowd counting rely on the powerful performance of deep learning models.
@@
 Subsequently, an accurate estimation of the total number of individuals is achieved by using a pixel-level crowd density image.
-To address the challenge of scale and density variations present in images, numerous methods have employed a multi-branch structure.
-This approach allows for the utilization of multi-scale features, effectively tackling the issue.
-Among them, some methods employ dilated convolutions with different dilation rates in a multi-branch structure, enabling the model to have receptive fields of different scales.
-Furthermore, to tackle challenges such as object occlusion, overlap in high density crowds, and the effects of lighting and background variations, almost all methods focus on enhancing the feature extraction capability of the model, primarily by selecting or designing various network architectures and employing different attention mechanisms.
+In order to tackle the issue of scale and density variations in images, many approaches have adopted a multi-branch structure.
+Among them, some methods employ dilated convolutions with different dilation rates in different branchs, enabling the model to have receptive fields of different scales.
+Furthermore, to tackle challenges such as object occlusion, overlap in high density crowds, and lighting and background variations, most methods focus on enhancing the feature extraction capability of the model, primarily by selecting or designing powerful network architectures.
 Optimizing the loss function yields significant effectiveness as well.
-For instance, in order to enhance model supervision, the loss function can be adjusted to ensure that the model establishes a more precise correspondence between the predicted density and the labeled location.
-Different research works adopt different perspectives and consider various rules.
-Through intricate and sophisticated network structural designs and algorithm optimizations, existing deep learning models have made remarkable progress in improving the accuracy of crowd counting.
-However, the complexity of the model to a large extent hinders subsequent researchers from modifying and further improving the model.
-For example, efforts to simplify existing models with complex structure to reduce the consumption of computing resources may significantly impact the accuracy of crowd counting and the models’ adaptability in complex scenarios.
-However, for structurally simple models, it is relatively easier to modify the models while minimizing the performance decrease.
-Hence, the reduction in complexity of the model structure is beneficial for the subsequent advancement of crowd counting.
-In the pursuit of simple structure but also need to consider the counting performance, achieving a balance between simplifying structure and enhancing performance is a challenge that we need to pay attention to.
-For the aforementioned reasons, our research objective is to design a crowd-counting model that strikes a balance between simplifying the model structure and improving counting performance for crowd counting.
-In terms of specific implementation, we first constructed a powerful backbone network, which is used to accurately capture the initial features of crowds from raw images.
-Next, we intricately designed a focus transition module, which efficiently integrates and refines key feature information in dynamic crowd scenes, ensuring the model’s effective adaptation to complex environmental changes.
-Finally, by implementing a practical and feasible multi-scale feature fusion strategy, we achieved deep integration and optimization processing of the extracted multi-level features.
-This ensures accurate identification and counting of crowds at different scales.
-Research shows that excellent crowd counting models can be designed with just a simple structure.
+For instance, to enhance model supervision, the loss function can be adjusted to ensure that the model establishes a more precise correspondence between the predicted density and the labeled location.
+In this paper, we focus on the impact of structural complexity on modifying models, discussing whether it is possible to find structurally simple models that achieve less performance decrease compared to complex models after modification.
+We strive to comprehensively consider the simplicity of the structure and the effectiveness of the model.
+Therefore, we opt for the ConvNeXT-Tiny architecture as our backbone, renowned for its simplicity and efficient design in convolutional neural networks.
+The multi-scale feature fusion structure we designed is very simple and allows us to effectively utilize information from different scales by fusing features from three different stages within the backbone.
+To simplify the structure, we only use concatenate operation on the features from three different branches.
+Moreover, we enhance the model’s effectiveness by including a carefully designed focus transition module before the fusion operation.
 In summary, this paper’s contributions can be outlined in three aspects:
-  1. 1.
-    We propose a crowd-counting neural network with a simple structure that consists of only a backbone and a multi-scale feature fusion structure, in contrast to other complex models.
-However, it achieves results comparable to the state-of-the-art, even with low parameters and computational requirements.
-  2. 2.
+  1. (1)
+    We present a crowd counting model with a simple structure that consists of only a backbone and a multi-scale feature fusion structure.
+Unlike other complex structure models, the fact that modifying our model, such as replacing different backbones, can achieve minimal performance loss demonstrates the effectiveness of this simple structure.
+Additionally, it achieves state-of-the-art (SOTA) results even with low parameters and computational requirements.
+  2. (2)
     We investigated three multi-scale feature fusion methods, including concatenate feature fusion, addition feature fusion, and stepwise addition feature fusion.
@@
 It efficiently extracts crowd features and effectively couples them across different scales, enhancing feature representation capability.
-  3. 3.
+  3. (3)
     To effectively extract key information and reduce dimensionality, we have developed a focus transition module that integrates dynamic convolution.
 It can dynamically adjust multi-dimensional weights based on feature attributes and optimize weights in channel and spatial dimensions.
-2 Related Work
-Crowd counting is a task of estimating the number of people in a specific location.
-Early research employed human detection methods [3, 4], where researchers used detectors to detect human in images and accumulated the count of each detected person to obtain the crowd count result in the image.
-The limitation of this method lies in the significant impact of occlusions between persons in crowded environments or densely packed crowds on the performance of detectors, leading to a decrease in counting accuracy.
-Subsequently, direct regression methods [5] were proposed, which count the crowd directly from images.
-However, this method often cannot provide detailed information about individual distribution and positioning, and it is sensitive to model assumptions and parameter selection, which can affect accuracy in complex or highly variable density scenes.
-Therefore, methods based on density map regression [6, 7] have become mainstream, where the final crowd count is obtained by integrating density map estimates from images.
-Relying on the powerful representation capability of convolutional neural networks (CNNs), the majority of methods in this field are built upon the framework of CNNs.
-There are also many methods leveraging the strong feature extraction ability of transformers.
-To further improve performance, researchers have developed and integrated various technological and architectural innovations.
-In this section, we will review the relevant work on crowd counting from three aspects: model design, loss function design, and algorithm strategy design.
-2.1 Model Design for Crowd Counting
-With the development of deep learning, various methods for model design in crowd counting have emerged.
-These methods include introducing multi-scale fusion strategies, integrating attention mechanisms, and employing transformers.
+Early research employed human detection methods [2], where researchers used detectors to identify humans in an image and then counted each detected person.
+The drawback of this method is that occlusions between individuals in crowded or crowded environments greatly affect the performance of detectors, resulting in a decrease in counting accuracy.
+Direct regression methods [3] directly estimate the number of people.
+However, this approach often lacks detailed information about individual distribution and positioning, and it is sensitive to model assumptions, which may impact accuracy in complex or highly variable density scenarios.
+Nowadays, density map regression methods [4] have become popular, as they integrate density map estimates from images to determine the crowd count.
+These methods rely heavily on strong representation capabilities of networks, with most of them utilizing convolutional neural networks (CNNs).
+Additionally, some methods also leverage transformers.
+To enhance performance, researchers have made advancements in technology and architecture.
+In this section, we review the relevant research on crowd counting, focusing on three aspects: model design, loss function design, and other aspects of the design.
+I-A Model Design for Crowd Counting
+Most of the research focuses on the following three aspects: utilizing powerful backbones, designing multi-scale feature fusion strategies, and integrating attention mechanisms.
+Choosing the right backbone is the primary task for building a crowd counting model.
+Given the remarkable success of transformer networks in the field of natural language processing (NLP), researchers have also begun to explore their application in crowd counting as a core feature extraction module.
+Sun et al.
+[5] utilize transformer models for feature extraction and information exchange by overlapping and concatenating image blocks and adding contextual tokens.
+Qian et al.
+[6] introduce a “U-shaped” multi-scale transformer network for crowd counting and introduce a new loss function to supervise regression, significantly improving counting accuracy.
 Multi-scale feature fusion methods attempt to address the challenge of large variations in crowd scale or size diversity by integrating features from multiple different levels.
@@
 [4] adopt a multi-column CNN architecture, where each column CNN utilizes filters with different receptive fields to adapt to scale variations.
-Jiang et al.
-[9] adopt an encoder-decoder structure that integrates multi-level features through skip connections across layers, aggregating low-level spatial details and high-level semantic information in a bottom-up manner.
-This enables accurate prediction from the entire image input to high-resolution density maps.
 Liu et al.
@@
 [8] introduce a multi-scale mechanism that utilizes information provided by different layers in the encoder to generate feature maps with varying receptive field sizes.
-Zhang et al.
-[12] designed a scaled adaptive convolutional neural network called SaCNN.
-This network combines features from multiple levels to enable the model to flexibly correspond to scale and viewpoint differences in the input image.
 Han et al.
 [9] employ a feature selection and inheritance learning approach, along with mask selection and inheritance loss, to perform multi-scale fusion.
+Wang et al.
+[10] use tree-based enhancers to address variations in crowd scale and leverages multi-level auxiliaries to filter out background pixels while adapting to changes in crowd density.
 These methods achieve good results by effectively integrating multi-scale features.
 Attention mechanisms are also an efficient method.
-Cao et al.
-[14] proposed SANet, which utilizes multi-scale convolutions and an adaptive global attention mechanism to handle crowd images of different sizes and densities.
 Jiang et al.
@@
 [12] employ a learnable global attention module and a local partition attention module, achieving decent performance in counting accuracy.
-Given the remarkable success of transformer networks in the field of natural language processing (NLP), researchers have also begun to explore their application in crowd counting as a core feature extraction module.
-Sun et al.
-[17] utilize transformer models for feature extraction and information exchange by overlapping and concatenating image blocks and adding contextual tokens.
-Qian et al.
-[18] introduce a “U-shaped” multi-scale transformer network for crowd counting and introduce a new loss function to supervise regression, significantly improving counting accuracy.
-In addition to the aforementioned methods, there are some other designs that contribute to crowd counting.
-The CSRNet proposed by Yan et al.
-[2] consists of a convolutional neural network serving as the frontend for 2D feature extraction, and it utilizes dilated convolutional neural networks as the backend to provide a larger receptive field and replace pooling operations.
-Similarly, Tian et al.
-[8], besides employing a transformer network that focuses more on spatial aspects, also utilize a regression head composed of dilated convolutions to expand the receptive field, achieving good counting results as well.
-2.2 Loss function design for crowd counting
-Many studies have designed loss functions tailored to the characteristics of crowd counting.
-Ma et al.
-[20] no longer rely on low-quality density maps generated from Gaussian kernels but construct a density contribution probability model from point annotations, using Bayesian inference for loss estimation.
 Wang et al.
-[21] treat crowd counting as a distribution matching problem, employing Optimal Transport (OT) to find the best match between the generated density map and the point annotations.
+[13] adopted a multi-layer context fusion approach and introduced a directive attention mechanism to improve accuracy.
+I-B Loss Function for Design Crowd Counting
+Some works have designed loss functions tailored to the characteristics of crowd counting tasks.
+Wang et al.
+[14] treated crowd counting as a distribution matching problem, employing Optimal Transport (OT) to find the best match between the generated density map and the point annotations.
 Yan et al.
-[22] proposed a multi-resolution loss function, PML, which can learn to better fit the posterior distribution using multi-scale information, thereby improving counting performance.
-2.3 Algorithmic Strategies design for crowd counting
-In addition to model design and loss function design, there are also studies that approach crowd counting from algorithmic perspectives.
+[15] proposed a multi-resolution loss function, progressive multi-resolution loss (PML), which can learn to better fit the posterior distribution using multi-scale information, thereby improving counting performance.
+I-C Other Aspects
+In addition to model and loss function design, there are also some studies improving crowd counting methods from other different aspects.
 Song et al.
@@
 Additionally, it employs a recursive rectangular window attention mechanism to improve computational efficiency.
-3 Method
-Figure 2: This figure presents the architectural designs of four different crowd counting models, including SASNet, STEERER, CCTrans, and our simplified architecture.
-With the continuous evolution and development of deep learning technology, high-performance crowd counting models have achieved significant improvements in accuracy, but they also face the problem of increasing model complexity.
-As depicted in Figure 2, the three exemplary high-performance models (CCTrans, STEERER, and SASNet) demonstrate exceptional accuracy, but their underlying structures are highly complex.
-Specifically, SASNet utilizes multiple decoder blocks along with a complex fusion header design and interleaved fusion mechanisms, resulting in a complex architecture.
-STEERER relies on the intricate and computationally demanding HRNet48 as its backbone network and designs a complex fusion module.
-Although the CCTrans appears to be lightweight, it actually adds more complexity to the model overall because it is based on the massive parameter Twins-SVT pre-training network, even though it only attaches a regression header after multi-branch fusion.
-Based on this, the core of our research is to conceive and build a more streamlined and efficient framework for crowd counting models.
-The fourth framework in Figure 2 is precisely the simplified solution we propose.
-This model adopts the compact yet powerful ConvNeXT-Tiny as the base backbone network, and it comprises a multi-scale feature fusion structure that includes a focus transition module.
-We refer to this model as Fuss-Free Network(FFNet).
-FFNet is specifically illustrated in Figure 3.
-Firstly, the input image goes through the feature extraction backbone in the framework to extract features at three different scales.
-Then, these three features pass through the focus transition module to generate more semantically rich features.
-Next, each stage’s features go through channel concatenation and spatially-aware fusion via the framework’s multi-scale fusion.
-Finally, regression heads are used to regress the density maps.
-The final density maps and the integrated counts will be used to construct the loss function.
+II Method
+Figure 2: This figure presents the architectural designs of four different crowd counting models, including SASNet, PET, CCTrans, and our simplified structure.
+In recent years, a series of high-performance crowd counting models based on deep learning have been proposed.
+The trade-off for achieving high performance in these models is an increase in the complexity of the model structure.
+Figure 2 illustrates a comparison between the network structures of three mainstream high-performance crowd counting models, namely CCTrans [18], PET [17], and SASNet [19], and our proposed model structure.
+More specifically, SASNet incorporates multiple decoder blocks, along with a sophisticated fusion header design and interleaved fusion mechanisms, resulting in a highly intricate architecture.
+PET employs the Quadtree structure to address the challenge of crowd counting and incorporates a complex Transformer-based encoder-decoder, substantially amplifying the model’s intricacy.
+Although the structure of CCTrans may seem simple, it actually introduces additional complexity to the overall model.
+This is because it is built upon the massive parameter Twins-SVT network, even though it only adds a regression header after multi-branch fusion.
+The fourth structure in Figure 2 illustrates the simplified solution proposed by us.
+We utilize the compact yet powerful ConvNeXT-Tiny as the base backbone and integrate only a multi-scale feature fusion structure with a focus transition module.
+This model is referred to as Fuss-Free Network (FFNet).
+As depicted in Figure 3, FFNet comprises solely of a backbone and a simplified multi-scale feature fusion module.
+The backbone is responsible for extracting features at three different scales.
+The simplified multi-scale feature fusion module includes three branches, each equipped with a focus transition module.
+The features at three different scales are individually processed through their corresponding focus transition modules, resulting in the generation of more semantically rich features specifically related to crowded scenes.
+Finally, the features from the three branches are concatenated together, and then processed through a 1×1 convolution layer to merge the features into a single channel, resulting in the final density map.
+The source code for this project is available on GitHub at https://github.com/erdongsanshi/FFNet.
 Figure 3: The network architecture of FFNet.
 $H$, $W$, $C$ denote the height, width, and number of channels of the features, respectively.
-Gap represents global average pooling, Mean denotes mean operation, Conv stands for convolution, $\times1$Conv denotes convolution with a $\times1$ kernel, Dyconv represents dynamic convolution, MLP refers to multilayer perceptron, Sigmoid is the activation function, and the circles with “×”, “+”, and “C” represent element-wise multiplication, element-wise addition, and channel concatenation operations, respectively.
-3.1 Feature Extraction Backbone
-Different from most crowd counting studies, we employ ConvNeXt-Tiny [24] as the feature extraction backbone for the multi-scale fusion framework.
-This network integrates key design principles of transformers, such as layer normalization and multi-head attention mechanisms, and implements them in the ConvNeXt block structure.
-This combination allows ConvNeXt-Tiny to not only maintain the powerful processing capability of CNNs for spatial local features but also significantly enhance the understanding and utilization of global information.
-Compared to the deep structures of VGG [25] and the residual learning of ResNet [26], ConvNeXt-Tiny provides a more efficient computational path and a simplified network architecture.
-Its streamlined and efficient characteristics ensure that the model can achieve superior performance in visual tasks such as image recognition compared to traditional models.
-Therefore, it also aligns better with our desire for a concise and efficient network design.
-For crowd counting tasks, dealing with objects of various scales has always been a significant challenge.
-A widely used solution is to generate features at different resolutions.
-Thus, we will utilize the ConvNeXt-Tiny pretrained model to build a multi-stage hierarchical feature extraction structure, where each stage can capture feature representations at different levels and resolutions.
-This approach enables us to better address the multi-scale challenges of crowd counting while ensuring the simplicity of the framework.
-3.2 Multi-scale Fusion Strategy
+And the circles with “×”, “+”, and “C” represent element-wise multiplication, element-wise addition, and channel concatenation operations, respectively.
+II-A Feature Extraction
+Given our research objective of developing a crowd counting model with a balanced approach to simplifying the model structure, we opt for ConvNeXT-Tiny [20] as the backbone due to its notable attributes of simplicity and efficiency.
+ConvNeXts series build upon the foundation of a standard ResNet [21] while drawing inspiration from newer models proposed after 2020, such as vision Transformer and ResNeXt [22], with their innovative design and training techniques, for instance, one can apply techniques such as adjusting the stage compute ratio, utilizing Inverted Bottleneck and Large Kernel Sizes, and substituting ReLU with GELU, among others.
+By incorporating these operations, ConvNeXt-Tiny is able to preserve the strong spatial processing capabilities of CNNs for local features while also greatly improving the comprehension and utilization of global information.
+ConvNeXt-Tiny is a smaller version of the original ConvNeXt model which provides a more efficient computational path and a simplified structure.
+Its streamlined and efficient design allows it to outperform traditional models in visual tasks like image recognition.
+This aligns with our goal of creating a concise and efficient network.
+In crowd counting tasks, dealing with objects of various sizes is a major challenge.
+One common approach is to generate features at different resolutions.
+To address this, To address this issue, we extract features from diverse stages of ConvNeXt-Tiny, thereby better adapting to the crowd counting task by capturing features at different hierarchical levels and resolutions from each stage.
+II-B Multi-scale Feature Fusion
 Figure 4: The three simple designs for multi-scale fusion are illustrated below.
 Given $S_{1}$, $S_{2}$, and $S_{3}$ as inputs, T represents the final output, and $T_{2}$ represents the output of the $S_{2}$ stage.
-For features at different levels of resolution, we need to align and fuse them.
-To avoid introducing too much additional computational cost and adhere to the principle of simplicity and efficiency, we propose three simple designs for multi-scale fusion.
-As shown in the figure 3, the “Concatenate” fusion directly applies transposed convolution to features $S_{3}$ and $S_{2}$ for upsampling to match the spatial size of $S_{1}$, and then fuses the features through channel concatenation.
+The main objective of multi-scale feature fusion is to enhance model performance by effectively integrating features at various scales, which enables the model to capture both fine-grained details and overall contextual information simultaneously.
+As a result, the model’s accuracy, robustness, and generalization capabilities are improved, allowing it to better handle variations in object size, pose, appearance, etc.
+In order to adhere to the principle of simplicity and efficiency, we evaluate three commonly used and simple multi-scale feature fusion strategies and select the most efficient one.
+As shown in the Figure 3, the “Concatenate” fusion directly applies transposed convolution to features $S_{2}$ and $S_{3}$ for upsampling to match the spatial size of $S_{1}$, and then fuses the features through channel concatenation.
 The “Addition” fusion first reduces the number of channels in features $S_{3}$ and $S_{2}$ to match the channel count of $S_{1}$.
@@
 The “Stepwise Addition” fusion involves progressively upsampling and combining features of different scales, similar to the design of FPN [23].
-In detail, these designs employ transpose convolution instead of interpolation operations because transpose convolution can better preserve and reconstruct the spatial structure information and feature hierarchy of the image.
-Compared to simple methods like bilinear interpolation for upsampling, transpose convolution enhances the output resolution by learning weight parameters, allowing for a more flexible and adaptive generation of high-quality feature maps.
-This is particularly significant for reconstructing complex crowd distribution scenes.
-The above three designs performed well in experiments, with the “Concatenate” fusion achieving the best results.
-We default to this design in the FFNet framework.
-3.3 Focus Transition Module
-Directly fusing features from different levels and resolutions of the backbone may lead to information redundancy due to misalignment in spatial dimensions and an excessive number of channels.
-This could potentially result in adverse effects such as decreased learning efficiency and inaccurate information propagation.
-Especially in crowd counting tasks, directly fusing features may fail to accurately retain spatial details from lower layers and lack an understanding of the global scene layout.
-This can lead to biases in estimating dense or sparse crowd distributions.
-Therefore, to address this issue and fully exploit the potential of the framework, we designed a focus transition module.
-It effectively focuses on dynamic and static features, achieving efficient dimensionality reduction and feature extraction to facilitate the transition.
-Inspired by [27], we employed a novel dynamic convolution in the design stage, as illustrated in Figure 5.
-Figure 5: Dyconv.
-FC stands for fully connected layer, Gap represents global average pooling, Conv denotes convolution, Sigmoid and Relu are activation functions, multiply refers to element-wise multiplication, $A_{nk}$, $A_{sk}$, $A_{ik}$, $A_{ok}$ represent attention along the dimensions of the number of convolution kernels, convolution kernel size, input channels, and output channels, respectively, and $W_{k}$ represents the weights of different convolution kernels.
-Its core idea is to dynamically generate convolution kernel weights during the model runtime based on the input data and the network’s learned state, rather than using predefined and fixed weight matrices.
-Dynamic convolution will flexibly capture features across the four key dimensions: input channels, output channels, kernel size, and the number of kernels within the convolution.
-It utilizes average pooling and convolution for attention and then dynamically weights them.
-It can adaptively adjust the parameters of the convolutional kernel to better match the features of the input data.
-This adaptive capability enables dynamic convolution to effectively extract features from data of different scales and complexities, thereby enhancing the model’s ability to handle diverse inputs.
-In terms of algorithms, we assume X to be the input of dynamic convolution.
-The process of dynamic convolution attention extraction and dynamic weighting can be defined as:
-$A(\mathbf{X})$ $=\sigma(\operatorname{Conv}(FC(\operatorname{AvgPool}(\textbf{X})))),$ (1)
-$Y(\mathbf{X})$ $=\mathbf{X}\odot\sum_{k=1}^{n}(\mathbf{A_{nk}}\odot\mathbf{A_{sk}}\odot\mathbf{A_{ik}}\odot\mathbf{A_{ok}}\odot\mathbf{W_{k}}),$ (2)
-where A(X) represents the attention calculation, AvgPool denotes average pooling, FC stands for fully connected layer, Conv indicates convolution operation, n represents the number of convolutional kernels, and $\sigma$ is the sigmoid activation function; Y(X) represents the dynamic convolution weighting process, $A_{nk}$, $A_{sk}$, $A_{ik}$, $A_{ok}$ are the attention weights on four dimensions: the number of convolution kernels, kernel size, input channels, and output channels, respectively, $W_{k}$ represents the weights of different convolution kernels, and C denotes the input optimized by weights.
-In terms of specific design, as illustrated in Figure 3, the entire structure combines feature weight optimization, dynamic perception, and residual components.
-The feature output from the backbone is first subjected to adaptive feature weight optimization.
-This mechanism strengthens and filters features along the channel dimension, ensuring that the model focuses on the features crucial for crowd statistics.
-Then, the stacked dynamic convolutions automatically adjust filter parameters for different inputs to capture dynamic and static feature changes in the scene more effectively, achieving feature extraction in the process.
-Finally, the module performs spatial feature weight optimization, allocating different attention to each pixel in the image, enabling the model to automatically focus on key areas of crowd distribution while ignoring irrelevant background noise.
-In addition, we embed residual structures throughout the entire process to ensure that the original low-level feature details are preserved and integrated with the higher-level abstract features refined through multiple convolutional layers.
-It is worth noting that while the focus transition module enhances key features, it also implicitly achieves feature dimensionality reduction.
-This dimensionality compression and optimization, while preserving the integrity of key information, help reduce the computational complexity of the model, thereby improving operational efficiency.
+For the purpose of reconstructing complex crowd distribution features, we specifically select transpose convolution instead of interpolation operations, as it offers notable advantages such as learning capability, flexibility, spatial accuracy, and the ability to handle irregular patterns.
+This choice allows for more precise and adaptive upsampling, making it particularly impactful in the context of complex crowd distribution scenes.
+After conducting experiments, we found that the above three fusion strategies showed satisfactory performance.
+Among them, the “Concatenate” fusion strategy yielded the best results.
+Therefore, we have chosen to adopt this design as the default fusion strategy in the FFNet framework.
+II-C Focus Transition Module
+The backbone was originally designed for classification tasks, so the features it outputs contain more general semantic information, which differs from the features required for crowd counting tasks.
+For example, the most crucial task in crowd counting is to extract features of the crowd.
+Although the backbone will adapt to crowd counting tasks during the training process, further processing after the three features at different scales will enhance the flexibility of the model, thus better adapting to crowd counting tasks.
+Furthermore, the direct fusion of the different scales features from backbone may result in information redundancy, potentially reducing learning efficiency.
+For the aforementioned reasons, we have developed a focus transition module that effectively focuses on dynamic and static features, enabling efficient dimensionality reduction and feature extraction to facilitate the transition.
+According to the aforementioned requirements, we have devised the focus transition module (FTM) to transform the features of the three branches prior to fusion, while also adhering to principles of simplicity and efficiency.
+Figure 3 illustrates the main components of FTM, comprising dynamic convolution based [24] feature transformation, a channel attention mechanism, and a spatial attention mechanism.
+To begin with, we utilize a channel attention mechanism to capture the interdependencies among different channels of the feature map extracted from the backbone network.
+This mechanism assigns weights to each channel according to their significance, enabling the branch to concentrate on pertinent features while suppressing less informative ones.
+Subsequently, the dynamic convolution based feature transformation component is employed to enhance the feature map by effectively capturing both dynamic and static feature variations in the scene.
+Lastly, we implement a spatial attention mechanism to capture the interdependencies within the spatial of the feature map.
+This mechanism assigns weights to various spatial locations, allowing the branch to focus on spatial regions while suppressing irrelevant or noisy ones.
+The most crucial component of FTM is the dynamic convolution based feature transformation, which includes dynamic convolution and residual connection.
+Dynamic convolution is employed in our model as an alternative to conventional convolution, effectively enhancing the model’s representational power without the need for increasing the depth or width of the network.
+Furthermore, we incorporate residual connections throughout the entire process to enhance gradient flow and facilitate the learning of intricate relationships.
+It is worth noting that while the FTM enhances key features, it also implicitly achieves feature dimensionality reduction.
+This dimensionality reduction, while preserving the integrity of key information, helpS reduce the computational complexity of the model.
 In detail, after each dynamic convolution, there is a ReLU activation layer and a BatchNorm layer to ensure the non-linear expression capability of features and the stability of normalization.
-The entire process of the focus transition can be represented by the following formula:
+The entire process of the FTM can be represented by the following formula:
 $C(\mathbf{F})$ $=\sigma(\operatorname{MLP}(\operatorname{AvgPool}(\mathbf{F})))\odot\mathbf{F},$ (1)
@@
 $S(\mathbf{P})$ $=\sigma(\operatorname{Conv}(\mathbf{\bar{P}}))\odot\mathbf{P},$ (3)
-where F is the input, C(·) is the channel feature weight optimization, Y(·) is the dynamic convolution, P(·) represents the output after stacked dynamic convolutions, S(·) is the spatial feature weight optimization, $\bar{P}$ is the mean value of P, and MLP is the multi-layer perceptron.
-The focus transition module we designed will fully unleash the potential of the streamlined framework, aiding the model in extracting richer, more targeted semantic information, achieving smooth transition effects, and enhancing the accuracy and robustness of crowd counting tasks.
-3.4 Loss Function
-The total loss consists of three components: count loss, optimal transport(OT) loss [20], and variation loss.
+where $\mathbf{F}$ is the input, $C(\cdot)$ is the channel feature weight optimization, $Y(\cdot)$ is the dynamic convolution, $P(\cdot)$ represents the output after stacked dynamic convolutions, $S(\cdot)$ is the spatial feature weight optimization, $\mathbf{\bar{P}}$ is the mean value of $\mathbf{P}$, and MLP is the multi-layer perceptron.
+The fundamental concept behind dynamic convolution is to dynamically combine multiple parallel convolution kernels based on their attentions, which are determined by the input data.
+This approach generates convolution kernel weights during the runtime of the model, allowing for adaptability to different input data.
+By capturing features across various dimensions such as input channels, output channels, kernel size, and the number of kernels, dynamic convolution offers flexibility in feature extraction.
+Dynamic convolution utilizes a combination of average pooling and convolution to determine attention weights, which are then dynamically adjusted.
+This allows for adaptive parameter adjustment of the convolutional kernel to better align with the input data’s features.
+We define $\mathbf{X}$ as the input of dynamic convolution and the process of dynamic convolution attention extraction and dynamic weighting as follows:
+$A(\mathbf{X})$ $=\sigma(\operatorname{Conv}(FC(\operatorname{AvgPool}(\textbf{X})))),$ (4)
+$Y(\mathbf{X})$ $=\mathbf{X}\odot\sum_{k=1}^{n}(\mathbf{A}_{nk}\odot\mathbf{A}_{sk}\odot\mathbf{A}_{ik}\odot\mathbf{A}_{ok}\odot\mathbf{W}_{k}),$ (5)
+where $A(\cdot)$ represents the attention calculation, AvgPool denotes average pooling, FC stands for fully connected layer, Conv indicates convolution operation, n represents the number of convolutional kernels, and $\sigma$ is the sigmoid activation function; $Y(\cdot)$ represents the dynamic convolution weighting process, $\mathbf{A}_{nk}$, $\mathbf{A}_{sk}$, $\mathbf{A}_{ik}$, $\mathbf{A}_{ok}$ are the attention weights on four dimensions: the number of convolution kernels, kernel size, input channels, and output channels, respectively, $\mathbf{W}_{k}$ represents the weights of different convolution kernels, and C denotes the input optimized by weights.
+II-D Loss Function
+We employed the Loss function utilized in DM-Count [14] which consists of three components: count loss, optimal transport (OT) loss [25], and variation loss.
 Count loss is a critical loss in crowd counting that quantifies the gap between the number of people predicted by the model and the actual number of people labeled, thus contributing to better convergence of the model.
-It is defined as the formula 6.
-Based on the counting loss, we further constructed a variation loss using the predicted density map and the ground truth density map.
-It employs the L1 norm to measure the intensity or grayscale variation differences between adjacent pixels in the predicted and ground truth density maps.
-Through this approach, our goal is to enhance the smoothness of the image and effectively reduce the adverse effects of potential noise on density estimation.
-Its definition is as follows:
-$\mathcal{L}_{c}$ $=\frac{1}{N}\sum_{i=1}^{N}\left|c_{p}-c_{g}\right|,\mathcal{L}_{v}=\frac{1}{N}\sum_{i=1}^{N}\left|x_{i}-y_{i}\right|,$ (6)
-where $c_{p}$ is the predicted population and $c_{g}$ is the actual population, N represents the batch size, $x_{i}$ represents the density value of the $i$-th pixel in the true density map, and $y_{i}$ represents the predicted density value of the $i$-th pixel at the corresponding position in the predicted density map.
+It is defined as follows:
+$\mathcal{L}_{c}(\mathbf{z},\hat{\mathbf{z}})$ $=\left|\|{{\mathbf{z}}}\|_{1}-\|{\hat{\mathbf{z}}}\|_{1}\right|,$ (6)
+where $\mathbf{z}$ represents the vectorized binary map for dot-annotation and $\hat{\mathbf{z}}$ is the vectorized predicted density map returned by network, $\|\cdot\|_{1}$ denotes the L1 norm.
 OTLoss is used to measure the difference between two probability distributions, making it particularly suitable for handling misaligned data distribution problems.
 It encourages the model to learn probability densities that are closer to the true distribution.
-It first calculates the distance matrix between the predicted density map and the true density map using the Euclidean distance.
-Then, it utilizes the Sinkhorn algorithm for iterative optimization, finding the optimal transportation plan from the source distribution to the target distribution, and obtaining the dual variables during this process.
+It utilizes the Sinkhorn algorithm [26] for iterative optimization, finding the optimal transportation plan from the source distribution to the target distribution, and obtaining the dual variables during this process.
 Finally, based on these dual variables and the predicted density distribution, OTLoss is computed.
-This loss measures the discrepancy between the predicted density distribution and the true labels in terms of optimal transportation.
-Throughout the entire process, the regularization parameter regulates the balance between distance cost and distribution smoothness.
-Its complete definition is as follows:
-$D_{ij}$ $=\sqrt{\left(x_{i}-x_{j}\right)^{2}+\left(y_{i}-y_{j}\right)^{2}},$ (7)
-where each matrix element $D_{ij}$ represents the Euclidean distance between the $i$-th grid point in the predicted density map, with coordinates ($x_{i}$, $y_{i}$), and the $j$-th ground-truth point in the true density map, with coordinates ($x_{i}$, $y_{j}$).
-$\mathbf{\beta}=\log(\mathbf{P})-\frac{1}{\lambda}\left[\log\left(\sum_{j=1}^{N}e^{-\lambda D_{ij}}\right)\right]_{i=1}^{M},$ (8)
-where $\beta$ represents the dual variables obtained through the Sinkhorn algorithm, $D$ is the distance matrix, $P$ represents the prediction map, $M$ represents the number of rows in matrix $D$, NN represents the number of columns, and $D_{ij}$ represents the value at the $i$-th row and $j$-th column of the distance matrix, and $\lambda$ is the regularization parameter set to 10.
-$\mathcal{L}_{ot}$ $=\sum_{i=1}^{N}\left[\left(\frac{\mathbf{\beta}_{i}}{P_{i}}-\frac{\mathbf{\beta}_{i}\cdot\mathbf{\rho}_{i}}{P_{i}^{2}}\right)\mathbf{\rho}_{i}\right],$ (9)
-where $N$ represents the batch size, $\beta_{i}$ is the dual variable computed by the Sinkhorn algorithm for the $i$-th sample, $P_{i}$ is the sum of predicted densities in the $i$-th sample, and $\rho_{i}$ is the vector composed of all predicted density values for the $i$-th sample.
-Finally, the total loss is the weighted sum of three components. Let $\lambda_{1}$ and $\lambda_{2}$ be the weighting values, the total loss is defined as:
-$\mathcal{L}_{all}$ $=\mathcal{L}_{c}+\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{2},$ (10)
-Combining counting loss, OTLoss, and variation loss can enhance the accuracy and robustness of crowd counting models.
-Additionally, regularization reduces noise in the predicted maps, improves model generalization, accelerates convergence, and helps train more accurate and stable crowd counting models.
-In our experiments, we fixed $\lambda_{1}$ at 0.1 and only fine-tuned $\lambda_{2}$.
-4 Experiment
-This section validates the effectiveness of our method on crowd counting benchmarks through extensive experiments and ablation studies, comparing it with state-of-the-art techniques.
-4.1 Implementation Details
-4.1.1 Dataset
-We evaluate our methods in four benchmarks, including UCF_CC_50 [29], ShanghaiTech Part A and B [6], and NWPU-Crowd [30].
-These datasets differ in terms of scene complexity, crowd density range, and image resolution.
-UCF_CC_50 [29] consists of 50 images of high-density crowds, with a total of 63,075 individuals across the entire dataset.
-The number of individuals in each image ranges from 94 to 4543, with an average of 1280 individuals per image.
-Extreme variations in crowd density and occlusion situations are among the challenges presented by the densely crowded complex scenes in this dataset.
-Due to the relatively small size of the dataset, Idrees et al.
-[29] defined a cross-validation protocol for training and validating models.
-ShanghaiTech [6] comprises a total of 1198 annotated images with 330,165 individually marked heads, divided into Part A and Part B.
-Part A consists of 482 images, among which 300 are allocated for training and 182 for testing; these images were sourced from the Internet.
-In contrast, Part B comprises 400 training images and 316 testing images, all sourced from real surveillance footage.
-These two parts of the dataset, due to their complex real-life scenarios, diverse environmental conditions, and distribution of crowds, pose significant challenges for crowd counting technology.
-They are crucial benchmarks for measuring and improving the accuracy of crowd counting algorithms.
-NWPU-Crowd [30] differs from the previous ones in several aspects.
-It has a larger scale, containing 5,109 high-resolution images covering various lighting conditions, perspectives, and scene types.
-The number of people ranges from tens to tens of thousands, providing more diverse scenes and a wider range of people counts.
-Additionally, the dataset includes a few negative samples that contribute to the model’s increased robustness.
-Compared to the previous datasets, it imposes higher precision requirements and tests the generalization ability of crowd counting algorithms.
-4.1.2 Training Settings and Hyperparameters
+Its definition is as follows:
+$\mathcal{L}_{ot}(\hat{\mathbf{z}})$ $=\langle\frac{\mathbf{\beta}^{\ast}}{\|{\hat{\mathbf{z}}}\|_{1}}-\frac{\langle\mathbf{\beta}^{\ast},\hat{\mathbf{z}}\rangle}{\|{\hat{\mathbf{z}}}\|^{2}_{1}},\hat{\mathbf{z}}\rangle,$ (7)
+[…]
+Furthermore, variation loss is utilized to measure the intensity or grayscale variation differences between adjacent pixels in the predicted and ground truth density maps.
+Specifically, the L1 norm is employed.
+We utilize this loss function to improve the overall image smoothness and effectively minimize the negative impact of potential noise on density estimation.
+To balance the contribution of losses from areas with different crowd densities, we use the actual crowd count as a weighting factor to adjust the variance loss.
+Its definition is as follows:
+$\mathcal{L}_{v}(\mathbf{z},\hat{\mathbf{z}})$ $=\|{{\mathbf{z}}}\|_{1}\cdot\||\mathbf{z}-\hat{\mathbf{z}}\|_{1}.$ (8)
+The total loss is the weighted sum of three components, which is defined as:
+$\mathcal{L}_{all}$ $=\mathcal{L}_{c}+\lambda_{1}\mathcal{L}_{ot}+\lambda_{2}\mathcal{L}_{v},$ (9)
+where $\lambda_{1}$ and $\lambda_{2}$ are the weighting values, and were assigned as 0.1 and 0.01, respectively, in the experiment.
+III Experiments
+III-A Implementation Details
+III-A1 Dataset
+We evaluate our methods in four benchmarks, including UCF_CC_50 [27], ShanghaiTech Part A (SHA) and ShanghaiTech Part B (SHB) [4], and NWPU-Crowd [28].
+The datasets, described in detail in TableI, vary in terms of scene complexity, crowd density range, and image resolution.
+Datasets Numbers of Image Average Resolution Total marked persons Max marked persons Min marked persons Average marked persons UCF_CC_50 50 2101 × 2888 63974 4633 94 1279 ShanghaiTech A(SHA) 482 589 × 868 241677 3139 33 501 ShanghaiTech B(SHB) 716 768 × 1024 88488 578 9 123 NWPU 5109 2191 × 3209 2133375 20033 0 418
+TABLE I: Crowd counting datasets used for evaluation.
+III-A2 Training Settings and Hyperparameters
 For a fair comparison, the backbone of FFNet is the official ConvNeXt-Tiny model, which is pretrained on the ImageNet-1k dataset.
 In our experiments, data augmentation was limited to random cropping and random horizontal flipping.
-For ShanghaiTech Part B, we set the cropping size to 512; for NWPU, the cropping size is set to 384; and for UCF_CC_50 and ShanghaiTech Part A datasets, we adjust the cropping size to 256.
+For SHB, we set the cropping size to 512; for NWPU, the cropping size is set to 384; and for UCF_CC_50 and SHA datasets, we adjust the cropping size to 256.
 The model training utilized the AdamW optimizer with a batch size of 8, and the initial learning rate was set to 1e-5.
@@
 All experiments were conducted using PyTorch on a single 16-GB RTX 4080.
-4.1.3 Evaluation Metrics
-In crowd counting research, Mean Absolute Error (MAE) and Mean Squared Error (MSE) are commonly used as evaluation metrics for crowd counting.
-The Mean Absolute Error (MAE) is the average of the absolute differences between the predicted counts and the actual counts.
-It provides an intuitive measure of how much the predicted values deviate from the actual values.
-The Mean Squared Error (MSE) is the average of the squared differences between the predicted counts and the actual counts.
-It is highly sensitive to outliers and requires the model to have good robustness to noise and outliers in the data.
-The expressions for MAE and MSE are as follows:
-$MAE$ $=\frac{1}{N}\sum_{i=1}^{N}\left|c_{i}-\hat{c}_{i}\right|,$ (11)
-$MSE$ $=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(c_{i}-\hat{c}_{i}\right)^{2}},$ (12)
-where N is the number of test images, and $c_{i}$ and $\hat{c}_{i}$ are the predicted and true number of people in the ith image, respectively.
-Method Venue Backbone Parameters FLOPS UCF_CC_50 SHA SHB NWPU(V) MAE MSE MAE MSE MAE MSE MAE MSE MCNN [6] CVPR16 CNN - - 377.6 509.1 110.2 173.2 26.4 41.3 218.5 700.6 CSRNet [2] CVPR18 VGG-16 - - 266.1 397.5 68.2 115.0 10.6 16.0 104.9 433.5 CAN [10] CVPR19 VGG-16 - - 212.2 243.7 62.3 100.0 7.8 12.2 93.5 489.9 SFCN [31] CVPR19 ResNet-101 - - 214.2 318.2 64.8 107.5 7.6 13.0 - - BL [20] ICCV19 VGG-19 - - 229.3 308.2 62.8 101.8 7.7 12.7 93.6 470.4 ASNet [15] CVPR20 VGG-16 - - 174.8 251.6 57.8 90.1 - - - - AMRNet [32] ECCV20 VGG-16 - - 184.0 265.8 61.5 98.3 7.0 11.0 - - DM-Count [21] NeurIPS20 VGG-19 - - 211.0 291.5 59.7 95.7 7.4 11.8 70.5 357.6 P2PNet [23] ICCV21 VGG-16 19.22M 104.71G 172.7 256.2 52.7 85.1 6.3 9.9 - - DKPNet [33] ICCV21 HRNet-W40 12.60M 46.31G - - 55.6 91.0 6.6 10.9 61.8 438.7 SASNet [38] AAAI21 VGG-16 38.90M 232.98G 161.4 234.46 53.5 88.3 6.3 9.9 - - MAN [16] CVPR22 VGG-19 30.96M 105.85G - - 56.8 90.3 - - - - CHFL [34] CVPR22 VGG-19 21.51M 108.27G - - 57.5 94.3 6.9 11.0 76.8 343.0 GauNet [35] CVPR22 ResNet-50 - - 186.3 256.5 54.8 89.1 6.2 9.9 - - CLTR [36] ECCV22 ResNet-50 40.95M 28.73G - - 56.9 95.2 6.5 10.6 61.9 246.3 FIDT [40] TMM22 HRNet-W48 66.58M 142.40G 171.4 233.1 57.0 103.4 6.9 11.8 51.4 107.6 PML [22] TCSVT23 ConvNeXt-S - - - - 53.4 93.1 6.2 9.7 - - LDFNet [37] arXiv23 VGG-19 - - 164.9 234.8 52.7 82.9 6.7 11.5 - - PET [19] ICCV23 VGG-16 51.65M 859.68G - - 49.3 78.9 6.2 9.7 58.5 238.0 STEERER [13] ICCV23 HRNet-W48 64.42M 94.40G - - 54.5 86.9 5.8 8.5 54.3 238.3 BCCT [17] arXiv21 T2T-ViT-14 - - - - 53.1 82.2 7.3 11.3 53.0 170.3 CCTrans [8] arXiv21 Twins-SVT 103.58M 99.44B 168.7 234.5 52.3 84.9 6.2 9.9 38.6 87.8 CUT [18] BMVC22 Twins-PCPVT 68.09M 90.45G - - 51.9 79.1 - - - - FFNet(ours) - ConvNeXt-T 29.02M 23.67G 161.1 233.3 48.3 80.5 6.1 9.0 41.2 113.2
-Table 1: Compared with the advanced methods on UCF_CC_50, ShanghaiTech Part A and B, and NWPU datasets.
-And compare the parameter quantity and flops with advanced methods.
-The best results are shown in bold, and the second-best results are underlined.
-4.2 Comparison with state-of-the-art methods
-4.2.1 Counting performance
+III-A3 Evaluation Metrics
+In crowd counting research, Mean Absolute Error (MAE) and Mean Squared Error (MSE) are commonly used as evaluation metrics for crowd counting. The MAE and MSE are defined as follows:
+$MAE$ $=\frac{1}{N}\sum_{i=1}^{N}\left|\mathbf{c}_{i}-\hat{\mathbf{c}}_{i}\right|,$ (10)
+$MSE$ $=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\mathbf{c}_{i}-\hat{\mathbf{c}}_{i}\right)^{2}},$ (11)
+where $N$ is the number of test images, and $\mathbf{c}_{i}$ and $\hat{\mathbf{c}}_{i}$ are the true and predicted number of people in the $i$-th image, respectively.
+Method Venue Backbone Parameters FLOPS UCF_CC_50 SHA SHB NWPU(V) MAE MSE MAE MSE MAE MSE MAE MSE CAN [7] CVPR19 VGG-16 - - 212.2 243.7 62.3 100.0 7.8 12.2 93.5 489.9 ASNet [11] CVPR20 VGG-16 - - 174.8 251.6 57.8 90.1 - - - - DensityCNN [29] TMM20 VGG-16 - - 244.6 341.8 63.1 106.3 9.1 16.3 - - DM-Count [14] NeurIPS20 VGG-19 - - 211.0 291.5 59.7 95.7 7.4 11.8 70.5 357.6 P2PNet [16] ICCV21 VGG-16 19.22M 104.71G 172.7 256.2 52.7 85.1 6.3 9.9 - - DKPNet [30] ICCV21 HRNet-W40 12.60M 46.31G - - 55.6 91.0 6.6 10.9 61.8 438.7 SASNet [19] AAAI21 VGG-16 38.90M 232.98G 161.4 234.46 53.5 88.3 6.3 9.9 - - PFDNet [31] TMM21 VGG-16 - - 205.8 289.3 53.8 89.2 6.5 10.7 - - MAN [12] CVPR22 VGG-19 30.96M 105.85G - - 56.8 90.3 - - - - CHFL [32] CVPR22 VGG-19 21.51M 108.27G - - 57.5 94.3 6.9 11.0 76.8 343.0 CLTR [33] ECCV22 ResNet-50 40.95M 28.73G - - 56.9 95.2 6.5 10.6 61.9 246.3 FIDT [34] TMM22 HRNet-W48 66.58M 142.40G 171.4 233.1 57.0 103.4 6.9 11.8 51.4 107.6 PML [15] TCSVT23 ConvNeXt-S - - - - 53.4 93.1 6.2 9.7 - - PET [17] ICCV23 VGG-16 51.65M 859.68G - - 49.3 78.9 6.2 9.7 58.5 238.0 STEERER [9] ICCV23 HRNet-W48 64.42M 94.40G - - 54.5 86.9 5.8 8.5 54.3 238.3 BCCT [5] arXiv21 T2T-ViT-14 - - - - 53.1 82.2 7.3 11.3 53.0 170.3 CCTrans [18] arXiv21 Twins-SVT 103.58M 99.44G 168.7 234.5 52.3 84.9 6.2 9.9 38.6 87.8 CUT [6] BMVC22 Twins-PCPVT 68.09M 90.45G - - 51.9 79.1 - - - - FFNet(ours) - ConvNeXt-T 29.02M 23.67G 161.1 233.3 48.3 80.5 6.1 9.0 41.2 113.2
+TABLE II: Compared with the advanced methods on four datasets. (The best results are shown in bold, and the second-best results are underlined).
+III-B Comparison with state-of-the-art methods
+III-B1 Counting performance
 CNN-based crowd counting methods dominate the field of crowd counting, with most approaches building upon the foundation of CNNs for model design.
-As shown in Table 1, our method outperforms them, exhibiting significant advantages in several benchmark tests and achieving state-of-the-art (SOTA) or near-state-of-the-art performance levels.
+As shown in Table II, our method outperforms them, exhibiting significant advantages in several benchmark tests and achieving SOTA or near-state-of-the-art performance levels.
 Even when compared to the popular transformer-based methods in recent years, our model demonstrates superiority, achieving better results in most cases.
-Specifically, we achieved outstanding results on the UCF_CC_50 dataset, with an MAE of 161.4, surpassing existing methods and reaching a new state-of-the-art level.
-On the ShanghaiTech Part A subset, we demonstrated significant superiority with an MAE value of 48.3, successfully surpassing the latest SOTA in this field.
-In the evaluation of the ShanghaiTech Part B subset, despite fierce competition, our method still achieved a highly competitive second-place performance with an MAE of 6.1.
+Specifically, we achieved outstanding results on the UCF_CC_50 dataset, with an MAE of 161.4, surpassing existing methods and reaching a new SOTA level.
+On the SHA subset, we demonstrated significant superiority with an MAE value of 48.3, successfully surpassing the latest SOTA in this field.
+In the evaluation of the SHB subset, despite fierce competition, our method still achieved a highly competitive second-place performance with an MAE of 6.1.
 Finally, in the more challenging NWPU-Crowd dataset, our method also achieved results close to the top tier.
-Without excessive complexity in design, our model achieves incredible results on these datasets.
-This fully demonstrates the effectiveness of FFNet.
-4.2.2 Parameters and Computational Resources
+Without excessive complexity in design, our model achieves excellent results on these datasets.
+III-B2 Parameters and Computational Resources
 Adhering to the original intention of our work, we aimed to simplify the model while achieving high crowd counting performance.
-Therefore, we conducted lightweight testing on FFNet and state-of-the-art methods.
+Therefore, we conducted lightweight testing on FFNet and SOTA methods.
 The testing primarily focused on open-source models released in the past three years.
@@
 Models with lower FLOPS consume fewer computational resources, have faster inference speeds, are more suitable for low-power environments, and are easier to deploy on mobile and embedded devices.
-The experimental testing results are presented in Table 1.
+The experimental testing results are presented in Table II.
 The data shows that over the past three years, as counting performance has continued to improve, there has been a significant increase in model parameters and FLOPs, resulting in greater model complexity.
@@
 While achieving superior counting performance, achieve optimal FLOPs and highly competitive model parameters.
-We believe that FFNet’s SOTA counting performance with such low FLOPs and model parameters is truly commendable.
 This effectively proves the efficacy of our proposed model: concise, efficient, low in complexity, and yet powerful in performance.
 FFNet effectively balances model simplification and performance enhancement, making it suitable not only for deployment in resource-constrained environments but also for performing exceptionally well in scenarios with high real-time requirements.
-This significantly drives the development and application of efficient crowd-counting technologies.
-4.3 The Effectiveness of Our Simplified Structure
-To validate the effectiveness of our concise network framework, we conducted corresponding experiments.
-We replaced FTM with ordinary convolutions.
-Specifically, we replaced FTM with four concatenated simple convolutions on top of a simple framework.
-Method SHA SHB
-MAE MSE MAE MSE
-SASNet [38] 53.5 88.3 6.3 9.9
-CCTrans [8] 52.3 84.9 6.2 9.9
-CUT [18] 51.9 79.1 - -
-PET [19] 49.3 78.9 6.2 9.7
-STEERER [13] 54.5 86.9 5.8 8.5
-FFNet+CNN 51.7 89.0 6.2 9.5
-FFNet+FTM 48.3 80.5 6.1 9.0
-Table 2: Simple network experiment: A comparison between a simple network with ordinary convolutions and a simple network with FTM against state-of-the-art methods.
-The results, as shown in Table 2, demonstrate that, compared to state-of-the-art approaches, our framework achieved an MAE of 51.7 on the ShanghaiTech Part A dataset and 6.2 on the Part B dataset, solely utilizing ordinary convolutions for intermediate downsampling and feature extraction.
-This surpasses most state-of-the-art methods.
-This further confirms the effectiveness of the streamlined network architecture we adopted and verifies the positive role of the FTM module in enhancing the overall network performance.
-4.4 Multi-scale fusion experiment
-In this section, we conducted comparative experiments on three simple multi-scale fusion strategies designed in Section 3.2.
-The experiment only adjusted the fusion method of the FFNet and validated it on the Part A and Part B datasets.
-The experimental results are shown in Table 3 below.
-In Part A, both the “Add” and “Stepwise addition” methods demonstrate good performance, but the “concatenate” method outperforms them.
-In Part B, the “concatenate” method shows better performance and robustness.
-This proves the effectiveness of the “concatenate” method in crowd counting tasks.
+III-C The Effectiveness of Our Simplified Structure
+To verify the effectiveness of our concise structure, we conducted replacement experiments on the backbone of the structure.
+We replaced the backbone of our model, CCTrans, and PET, with different pre-trained networks.
+Subsequently, we performed parameter fine-tuning on the modified models mentioned above using the SHA dataset, and conducted a count of the number of parameters and Flops for each model after replacing the backbone, using a 512×512 input.
+The results are shown in Table III.
+When using ConvNeXt-Tiny and Swin-Transformer-Small, our structure achieved significantly lower MAE and MSE with a minimal number of parameters and floating-point operations compared to the other CCTrans and PET.
+In addition, when using MobileNet-v3 as the backbone, our structure achieved MAE and MSE that are second only to CCTrans, with the lowest number of parameters (3.05M) and floating-point operations (1.51G).
+The experimental results show that our structure exhibits better robustness and effectiveness with lower parameter and computational complexity compared to other the structures of CCTrans and PET.
+Method MobileNet-v3 ConvNeXt-Tiny Swin-Transformer-Small Parameters FLOPS MAE MSE Parameters FLOPS MAE MSE Parameters FLOPS MAE MSE CCTrans [18] 4.74M 9.08G 62.9 102.1 32.12M 41.01G 53.9 86.5 37.35M 47.78G 56.6 92.4 PET [17] 15.19M 133.81G 66.6 110.6 90.96M 343.54G 54.1 85.1 106.68M 404.46G 58.1 90.0 FFNet 3.05M 1.51G 64.3 109.3 29.02M 23.67G 48.3 80.5 34.26M 30.44G 54.7 88.1
+TABLE III: Backbone replacement experiment. (Comparision of the effectiveness of different structs.)
+We also compare our model with recent lightweight crowd counting models using MobileNet-v3 as the backbone.
+As show in Table IV, by simply using a lightweight backbone, our model easily achieves lightweight performance.
+Method Parameters FLOPS(input size) SHA
+MAE MSE
+ECCNAS-Lat [35] 3.79M 7.38G(512×512) 66.3 116.5
+ECCNAS-Acc [35] 3.88M 16.46G(512×512) 62.0 110.9
+LMSFFNet [36] 4.58M 14.9G(1024×768) 85.8 139.96
+FFNet 3.05M 1.51G(512×512) 64.3 109.3
+(MobileNet-v3) 4.53G(1024×768)
+TABLE IV: Comparison with advanced lightweight models. (We utilized 256×256 cropped images to calculate the MAE and MSE metrics for the models.)
+III-D Multi-scale feature fusion
+In this section, we conducted comparative experiments on three simple multi-scale fusion strategies in Figure 4.
+The experiment only adjusted the fusion method of the FFNet and validated it on the SHA and SHB datasets.
+The experimental results are shown in Table V below.
+In SHA, both the “Addition” and “Stepwise addition” operations demonstrate good performance, but the “Concatenate” operation outperforms them.
+In SHB, the “Concatenate” operation shows better performance and robustness, which proves the effectiveness of the “Concatenate” operation in crowd counting tasks.
 Fusion Strategy SHA SHB
 MAE MSE MAE MSE
-Add 52.6 87.5 6.4 10.8
+Addition 52.6 87.5 6.4 10.8
 Stepwise addition 52.3 91.9 7.0 11.7
 Concatenate 48.3 80.5 6.1 9.0
-Table 3: Results of three simple multi-scale fusion designs on ShanghaiTech Part A and B datasets.
-4.5 Ablation Study
+TABLE V: Comparison of three simple multi-scale fusion designs on SHA and B datasets.
+III-E Ablation Study
 As mentioned earlier, our proposed FFNet consists of the following components: a feature extraction backbone, focus transition modules, and a multi-scale fusion structure.
-We conducted extensive ablation experiments on the ShanghaiTechA dataset to validate the effectiveness of each component in FFNet.
-4.5.1 Focus Transition Module
-We conducted two ablation experiments on the FTM module, one focusing on local effects and the other on global effects, to verify its effectiveness.
-First, at the local level, we tested the counting results of the three branches of the model before simple fusion on dataset A.
-Then, based on this, we removed the FTM module and tested the counting results of the three branches again, and compared the results.
-The effectiveness of FTM will demonstrate by observing the difference of counting performance before and after the three branches.
-At the global level, we tested the performance of FFNet with and without FTM to highlight the overall contribution of FTM to FFNet.
-The results are shown in Table 4.
-The results indicate that employing the FTM module, whether in local branches or in the overall model, effectively improves the counting accuracy, demonstrating the practicality of the focus transition module.
+We conducted extensive ablation experiments on the SHA dataset to validate the effectiveness of each component in FFNet.
+III-E1 Focus Transition Module
 Method SHA
@@
 […]
-Table 4: Effective improvement of model counting performance by FTM.
-4.5.2 Loss Function
-We also quantitatively evaluated the weighted sum composed of counting loss, variation loss, and optimal transport loss to verify the specific contributions of each loss function to the model’s performance.
-The results are shown in Table 5.
-Experimental results show that the model performance reaches its optimal state when the $\lambda_{1}$ and $\lambda_{2}$ are 0.1 and 0.01, respectively.
-This weighting scheme highlights the central role of count loss in crowd counting prediction, while also moderately considering the image smoothness brought by variation loss and the fine depiction of complex scenes by OT Loss.
-Loss function SHA
-MAE MSE
-$L_{c}$ 55.1 95.3
-$L_{c}+L_{ot}+L_{v}$ 55.3 94.9
-$L_{c}+\lambda_{1}L_{ot}+\lambda_{2}L_{v}(\lambda_{2}=0.1)$ 54.6 90.4
-$L_{c}+\lambda_{1}L_{ot}+\lambda_{2}L_{v}(\lambda_{2}=0.01)$ 48.3 80.5
-Table 5: The total loss function achieved optimal performance when weighting three loss functions and setting $\lambda 1$ and $\lambda 2$ to 0.1 and 0.01, respectively.
-4.5.3 Composition Analysis of FFNet
-Table 6 displays the component analysis of our model. For the ShanghaiTech PartA dataset, both FTM and the final “Concatenate” fusion contribute significantly to the final MAE metric.
+TABLE VI: Effective improvement of crowd counting performance through FTM.
+We conducted two ablation experiments on the FTM module, one focusing on local effects and the other on global effects, to verify its effectiveness.
+First, at the local level, we tested the counting results of the three branches of the model before simple fusion on dataset A.
+Then, based on this, we removed the FTM module and tested the counting results of the three branches again, and compared the results.
+The effectiveness of FTM will demonstrate by observing the difference of counting performance before and after the three branches.
+At the global level, we tested the performance of FFNet with and without FTM to highlight the overall contribution of FTM to FFNet.
+The results are shown in Table VI.
+The results indicate that employing the FTM module, whether in local branches or in the overall model, effectively improves the counting accuracy, demonstrating the practicality of the focus transition module.
+III-E2 Component Analysis of FFNet
 Method SHA
@@
 backbone+FTM+Concatenate 48.3 80.5
-Table 6: Composition analysis of FFNet.
-4.6 Visualization
-In order to further explore the internal mechanism of the model, we visualized the effective receptive field maps and their heat maps of the three branches of the model before simple fusion, so as to reveal the spatial coverage and response intensity distribution of the input data features of each branch of the model before simple fusion.
-The effective receptive field [39] refers to a pixel on the output feature map of a certain layer in a neural network whose response is influenced by the size of the input image, reflecting the actual sampling range of the layer on the input spatial size.
+TABLE VII: Component analysis of of FFNet.
+Table VII displays the component analysis of our model. For the SHA dataset, both FTM and the final “Concatenate” fusion contribute significantly to the final MAE metric.
+III-F Effective Receptive Field and Heatmap
+Figure 5: Three branches of the model pass FTM module in the receptive field map and their heat map.
+To enhance the intuitive understanding of our model’s internal mechanism, we visualized the effective receptive field (ERF) [37] as well as the heat maps of the three branches before fusion.
+This allows us to uncover the spatial coverage and response intensity distribution of the input data features in each branch before the fusion occurs.
+The theoretical Receptive Field (RF) denotes the area in the input space that directly influences the output of an individual neuron in a neural network.
+However, the contribution of each pixel within a receptive field to the response of an output unit may not be equal.
+The effective receptive field is employed to quantify the extent to which each pixel on the input feature map affects a pixel on the output feature map.
+Typically, the impact within a receptive field follows a Gaussian distribution.
 The heatmap is a visualization tool that represents numerical intensity through color coding and is used to display the model’s level of attention or confidence distribution towards specific targets or features in the input data.
@@
 The color distribution of the heatmaps also becomes more concentrated, demonstrating the model’s ability to integrate global information and recognize the overall distribution and density of the crowd.
-Figure 6: Three branches of the model pass FTM module in the receptive field map and their heat map
+Figure 6: FFNet visualizations on the most challenging NWPR populations.
+Six images from different scenes, with different scales, densities, and lighting differences, show the counting and fitting capabilities of FFNet.
 Finally, we visualize the model outputs using six images in the NWPU dataset, which exhibit varying scales, densities, lighting conditions and from different scenes.
 The visualization of the model’s final outputs is presented in Figure 6.
-Our model performs excellently on images of various scales, demonstrating good technical accuracy.
-Figure 7: FFNet visualizations on the most challenging NWPR populations.
-Six images from different scenes, with different scales, densities, and lighting differences, show the counting and fitting capabilities of FFNet.
-5 Conclusion
-In this paper, we propose a concise and efficient crowd counting model called Fuss-Free Net work (FFNet), which consists only of a backbone network, a multi-scale fusion structure, and focus transition modules.
-In the concrete implementation, we use the ConvNeXt-Tiny network to build a multi-branch structure, each branch is equipped with a focus transition module, and then the multi-branch features are fused through a simple cascading fusion strategy.
-Finally, experiments show that the counting accuracy of the model is equal to or even better than that of the complex model, despite its compact structure, small number of parameters, and low computational complexity, breaking the inherent relationship between complexity and performance improvement.
-FFNet provides a simple and efficient solution for crowd-counting tasks, but there is still plenty of room to explore.
+Our model performs excellently on images of various scales, demonstrating good the fitting ability.
+IV Conclusion
+In this paper, we propose a simplified and efficient crowd counting model called Fuss-Free Network (FFNet), which consists only of a backbone network, a multi-scale fusion structure with focus transition modules.
+We use the ConvNeXt-Tiny as backbone network and build a simple feature fusion that contains three branches.
+Each branch is equipped with its focus transition module, which has the capability to enhance the model’s flexibility and extract a greater variety of features related to the crowd.
+The features from three branches are fused through a concatenate operation and a $1\times 1$ convolution is applied to generate a single-channel density map.
+The experiment results show that the counting accuracy of the model is equal to or even better than that of the complex models, despite its compact structure, small number of parameters, and low computational complexity.
+FFNet provides a simple and efficient solution for crowd counting tasks, but there is still plenty of room to explore.
 For example, cross-domain adaptability, interpretability, and robustness are improved.